### v1

# Quantum phase estimation based filtering: performance analysis and application to low-energy spectral calculation

Abstract
Filtering is an important technique in quantum computing used for isolating or enhancing some specific states of quantum many-body systems. In this paper, we analyze the performance of filters based on the quantum phase estimation (QPE) algorithm, in which filtering removes states associated with bitstrings in the ancilla register above a given threshold. We show that when the conventional rectangular window function is used for the QPE input state, the resulting filter function exhibits an oscillating behavior known as the Gibbs phenomenon. We also show that in the case of the sine and Kaiser windows, this phenomenon is suppressed. Furthermore, we perform numerical simulations to compare the number of necessary queries to the Hamiltonian time evolution operation of for the QPE-based filtering algorithm and the quantum eigenvalue transformation of unitary matrices with real polynomials (QETU). We find that the number of queries required for Kaiser window-based filtering is comparable to that for QETU with optimized phase angles. As an application of the QPE-based filter, we also study a two-step algorithm for low-energy spectral simulations, composed of a coarse grid for filtering and a fine grid for obtaining final high-resolution spectra. As a benchmark of the proposed scheme for realistic continuous spectra, we present the density-of-states (DOS) calculation of antiferromagnetic type-II MnO in a one-particle approximation.
I introduction
Calculation of the electronic structure of complex materials is one of the most promising industrial applications of quantum computing. The last decade has seen rapid development of quantum-classical hybrid approaches that leverage currently available noisy intermediate-scale quantum computers, such as the variational quantum eigensolver [1] and recently-proposed sampling-based approaches [2, 3]. In addition, with the development of quantum hardware and quantum error correction techniques, there has also been remarkable progress on the development of algorithms designed for fault-tolerant quantum computers [4], such as quantum phase estimation (QPE) [5]. These fault-tolerant quantum algorithms are expected to yield true quantum advantage, that is, exponential speedup over classical algorithms.
Many industrially and fundamentally important phenomena in chemistry and materials science occur at a small energy scale involving the electronic ground state and low-lying excited states of a material. For example, the band gaps of typical semiconductors are of the order of $1$ eV, and the energy scale of many relevant types of chemical reactions, high-temperature superconductivity, and spintronics are in meV range. On the other hand, the total energy and the L2 norm of the Hamiltonian $\|H\|_{2}$ of materials scale with the size of the system and even for small molecules exceed $10-100$ eV. As the Hilbert space of the Hamiltonian increases, the ratio of the energy region of interest to $\|H\|_{2}$ becomes smaller. As a result, it becomes increasingly difficult to obtain low-energy spectra of the system with existing quantum algorithms, as observed in recent sampling-based quantum-classical hybrid calculations [3, 6].
To overcome this fundamental problem, quantum filtering algorithms have been developed to obtain the desired quantum state from a given input state. Examples of the ground-state filtering include the cosine [7], Gaussian [8, 9], and Lorentzian [10]-based filtering schemes. Quantum signal processing (QSP) is a versatile tool to encode arbitrary polynomial functions on a quantum circuit with a set of single-qubit rotation angles determined to construct a specified polynomial function, which offers another way of filtering [11]. Quantum eigenvalue transformation of unitary matrices with real polynomials (QETU) [12] is proposed as a simple QSP-based filtering technique with only one ancilla qubit, using controlled Hamiltonian evolution as a building block.
Utilizing QPE is yet another way of filtering quantum states. In early work, Poulin and Wocjan [13] proposed a ground-state preparation scheme based on a repeated backward operation of QPE. A separate systematic approach of filtering general energy intervals has been proposed in Ref. [14]. More recently, in Ref. [15] it is reported that the QPE-based approach with a coarse frequency grid filters out the excited states of a simple model system more efficiently than QETU. The QPE-based approach requires more ancilla qubits than QETU, but has an advantage that no procedure for finding QSP angles is needed.
Here we focus on the QPE-based filtering approach, combined with the window function techniques for QPE. Recently, the window function techniques have been attracting much interest in QPE. We perform a detailed analysis on how the quality of the QPE-based filter depends on the three window functions: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. The rectangular window, which is used in the conventional QPE algorithm, is known to suffer from the spectral leakage problem [16, 17]. The sine window is known to mitigate the problem and has been applied in quantum metrology [18, 19, 20, 21] to achieve the so-called Heisenberg-limit scaling. The Kaiser window [22] is proposed as an approximation of the prolate spheroidal window or the Slepian window [23, 24, 25, 26], and recently advantages of using this window over QSP-based approaches have been reported [27, 28].
Our main focus is not to increase the probability of obtaining the ground state of the system, as done in many ground-state preparation techniques, but to increase the probability of the states in the specified small energy region without altering their relative amplitudes. This requirement is important when calculating the low-energy spectral properties of the system.
We show that the rectangular window-based filter exhibits an oscillation known as the Gibbs phenomenon, and the oscillation is mitigated when the sine and Kaiser windows are used. The error in the Kaiser window-based filter decreases exponentially with the QPE frequency grid points, which leads to the same $O(\delta^{-1}\log\epsilon^{-1})$ scaling as the QETU-based filter. Here $\delta$ and $\epsilon$ are the width of the transitional region and the error parameter, respectively. We also numerically demonstrate that the Kaiser window-based filter and QETU with phase angles determined via convex optimization require a comparable number of Hamiltonian evolution operations for a given set of ($\delta,\epsilon$), indicating that the Kaiser window-based filter can be used as a potential alternative to the QETU-based filter. Finally, we consider an application of the proposed filtering approach to the QPE-based calculation of spectral functions [29, 30, 31, 32, 33, 34, 35]. Specifically, we consider a two-step QPE approach similar to the one proposed in Ref. [15], that uses a coarse QPE grid for filtering and a fine grid for obtaining high-resolution spectra. We test the present scheme by calculating the density-of-states (DOS) of MnO in a one-particle approximation [36] as an example of realistic spectra.
II theory
II.1 General formalism
Figure 1 shows the general form of the quantum circuit for QPE, which consists of $n$ and $s$ qubits for the ancilla and system registers, respectively.
Figure 1: Quantum circuit for QPE.
Here $U_{\textrm{in}}$ constructs the input state for the ancilla register
$U_{\textrm{in}}|0\rangle^{\otimes n}=\sum_{j=0}^{N-1}a_{j}|j\rangle$ (1)
with $N=2^{n}$. The real coefficients $a_{j}$ are related to the window function in classical signal processing, and specific forms of $a_{j}$ are discussed in the next session.
The state-preparation operation $U_{\psi}$ acts on the system register as
$U_{\psi}|0\rangle^{\otimes s}$ $=$ $|\psi\rangle$ (2)
$=$ $\sum_{\mu}C_{\mu}|\phi_{\mu}\rangle,$
where $\phi_{\mu}$ are the eigenstates of the Hamiltonian $H$ with eigenenergies $E_{\mu}$, and $C_{\mu}$ are expansion coefficients normalized as $\sum_{\mu}|C_{\mu}|^{2}=1$. After the state-preparation step, the controlled time evolution $U_{HT}=\textrm{exp}[iHT]$ is performed, where $T$ is a time duration chosen to scale the energy range of the system in $[0,2\pi]$. Throughout this work $\hbar=1$. Finally, the inverse of quantum Fourier transform (QFT) is performed. The quantum state before the measurement in Fig. 1 can be represented as
$|\Psi\rangle$ $=$ $\sum_{y=0}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle,$ (3)
where
$A_{\mu}(y)=\frac{1}{\sqrt{N}}\sum_{j=0}^{N-1}a_{j}\textrm{exp}[i(E_{\mu}T-\frac{2\pi}{N}y)j].$ (4)
The probability of measuring $y=0,1,\dots,N-1$ for each state $|\phi_{\mu}\rangle$ is
$P_{\mu}(y)=|A_{\mu}(y)|^{2},$ (5)
and the total probability of measuring $y$ is given as the sum over $\mu$,
$P(y)=\sum_{\mu}|C_{\mu}|^{2}P_{\mu}(y).$ (6)
For $N\to\infty$, each $P_{\mu}(y)$ in Eq. (5) has a peak at $E_{\mu}$ when seen as a function of frequency points
$\omega_{y}=\frac{2\pi}{NT}y.$ (7)
However, since $P_{\mu}(y)$ is defined only on the discrete frequency points for a finite $N$, it has, in general, nonzero tail probabilities, a phenomenon known as spectral leakage.
The QPE-based filtering amounts to discarding states with $y>y_{c}$ or $\omega_{y}>\omega_{y_{c}}$ in Eq. (3), where $y_{c}$ is a cutoff which is taken as a positive integer. To see the concept clearly, we rewrite the state in Eq. (3) as
$|\Psi\rangle$ $=$ $\sum_{y=0}^{y_{c}}\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}\Bigl{[}\frac{A_{\mu}(y)}{\sqrt{\mathcal{R}_{\mu}}}|y\rangle\Bigr{]}|\phi_{\mu}\rangle$ (8)
$+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle$
$=$ $\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}|\chi_{\mu}\rangle|\phi_{\mu}\rangle$
$+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle.$
Here, $\mathcal{R}_{\mu}$ is the renormalization factor, which depends on the state $|\phi_{\mu}\rangle$ through $P_{\mu}(y)$ in Eq. (5) as
$\mathcal{R}_{\mu}=\sum_{y=0}^{y_{c}}P_{\mu}(y),$ (9)
and
$|\chi_{\mu}\rangle=\sum_{y=0}^{y_{c}}\frac{A_{\mu}(y)}{\sqrt{\mathcal{R}_{\mu}}}|y\rangle$ (10)
are normalized $\mu$-dependent states. Thus, the QPE-based filtering amounts to extracting the first term of Eq. (8). One way to do this truncation is a coherent (i.e., fully quantum) filtering without measurement, by first performing an ideal quantum amplitude amplification (QAA) [37, 38, 39] and then tracing out the ancilla register. Another approach is through classical post-selection of the ancilla states after the measurement, as depicted in Fig. 1. In this case, the result becomes probabilistic, and the success probability can be increased by performing QAA on the corresponding states. It should be noted that that the resulting outcome in this classical filtering approach is a probability distribution, rather than a quantum state (the first term of Eq. (8)). This approach therefore differs from usual quantum filters but can be applied to enhance the spectral estimation, as discussed in Section III.
Figure 2: Errors in the QPE-based filtering. The filled regions indicate errors due to the truncation. (a) $P_{\mu}(y)$ for states with $E_{\mu}\leq\omega_{y_{c}}$. (b) $P_{\mu}(y)$ for states with $E_{\mu}>\omega_{y_{c}}$.
In both quantum and classical filtering approaches, the net effect of filtering is represented as the renormalization of the weight of each state $|\phi_{\mu}\rangle$ (see the Appendix for additional explanation on the classical approach):
$|C_{\mu}|^{2}\to|C_{\mu}|^{2}\mathcal{R}_{\mu}.$ (11)
The ideal filtering corresponds to $\mathcal{R}_{\mu,\textrm{ideal}}=1$ for the states with $0\leq E_{\mu}\leq\omega_{y_{c}}$ and $\mathcal{R}_{\mu,\textrm{ideal}}=0$ otherwise. The deviation from this ideal filtering, denoted as $\Delta\mathcal{R}_{\mu}$, is given as
$\Delta\mathcal{R}_{\mu}=\Bigl{\{}\begin{array}[]{ll}\sum_{y=y_{c}+1}^{N-1}P_{\mu}(y)&E_{\mu}\leq\omega_{y_{c}}\\ \sum_{y=0}^{y_{c}}P_{\mu}(y)&E_{\mu}>\omega_{y_{c}},\end{array}\Bigr{.}$ (14)
where $\sum_{y=0}^{N-1}P_{\mu}(y)=1$ is used. These errors originate from the tail part of the probability distribution due to spectral leakage, as schematically shown in Fig. 2. Here the errors are indicated by the filled regions in Figs. 2(a) and 2(b) for states with $E_{\mu}\leq\omega_{y_{c}}$ and $E_{\mu}>\omega_{y_{c}}$, respectively. Note that due to periodicity the QPE-based filtering has two boundaries, at $y=y_{c}$ and at $y=0$.
Although it is not an essential point of this approach, in this work we choose $y_{c}$ as $y_{c}=2^{n-m}-1$ with $m(<n)$ a positive integer. With this choice, $\omega_{y_{c}}=\frac{2\pi}{NT}y_{c}\approx\frac{2\pi}{T}\frac{1}{2^{m}}$, and we post-select or amplify the states whose first $m$ bits in the ancilla register are 00 (ex. $|{\bf{00}}1101\rangle$ when $m=2$). This filtering corresponds to a low-pass filter, but it is straightforward to consider filtering for general frequency intervals [14].
II.2 Window functions
Figure 3: (a) $a_{j}$ in Eq. (1) for the three windows. (b) $P_{\mu}(y)$ for the three windows and (c) $P_{\mu}(y)$ in the logarithmic scale. The vertical dotted lines in (b) and (c) indicate the position of $E_{\mu}$. Here $N=64$.
Figure 4: Error estimate in the Kaiser window-based filter defined in Eq. (23) for $n=6$ ($N=64$) and $n=10$ ($N=1024$). The horizontal dashed line indicates $\epsilon_{\textrm{max}}=10^{-7}$ used in Sec. II.5.
The performance of the QPE-based filtering depends crucially on the form of $a_{j}$. In this work, we consider three forms of $a_{j}$: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window.
The first window is the rectangular window characterized by
$a_{j}=\frac{1}{\sqrt{N}},$ (15)
which is used in the conventional QPE algorithm and corresponds to setting $U_{\textrm{in}}=\textrm{H}^{\otimes n}$ with H the Hadamard gate. The root mean squared error (RMSE) of the estimated phase achieved by this window function is proportional to $1/\sqrt{N}$ known as the standard quantum limit.
The second window is the sine window originally proposed in quantum metrology [18, 19, 20, 21]:
$a_{j}=\sqrt{\frac{2}{N}}\sin\frac{\pi}{N}j.$ (16)
In this case, the resulting RMSE is proportional to $1/N$ [20], which is known as the Heisenberg-limited scaling. The corresponding $U_{\textrm{in}}$ can be constructed in $O(n)$ complexity [40].
The third window is the Kaiser window [22, 41, 27, 28], which is introduced as an approximation of the prolate spheroidal window [23, 24, 25, 26]. We consider the following form:
$a_{j}\propto I_{0}\left(\pi\alpha\sqrt{1-\left(\frac{2j}{N}-1\right)^{2}}\right),$ (17)
where $I_{0}$ is the modified Bessel function of the order zero and $\alpha$ is a positive parameter. Note that our form is slightly different from other works [41, 27, 28], in that our definition $j$ goes from 00 to $N-1$. The Kaiser window and the prolate spheroidal window are more difficult to implement than the first two, but there have been proposals [42, 26]. In particular, the method of Ref. [42] uses quantum eigenvalue transformation with $O(n\sqrt[4]{\alpha+1}(\alpha+\log\epsilon^{-1}))$ gate complexity within error $\epsilon$.
We compare the results of all three windows in Fig. 3. In Fig. 3(a) we plot $a_{j}$ for the three windows. The Kaiser window has a Gaussian-like structure more localized than the sine window, and its spread is controlled by the parameter $\alpha$. In Figs. 3(b) and 3(c) we compare $P_{\mu}(y)$ defined in Eq. (5) for the three windows, for a specific example of $|\phi_{\mu}\rangle$ with $\frac{NT}{2\pi}E_{\mu}=10.2$ and $N=64$. The $P_{\mu}(y)$ in the rectangular window has the highest peak but has large leakage, and the sine window result shows a smaller tail probability [31]. For these two windows, the expression of $A_{\mu}(y)$ in Eq. (4) can easily be obtained; for the rectangular window
$A_{\mu}(y)$ $=$ $\frac{1}{N}\frac{1-e^{iE_{\mu}TN}}{1-e^{i\Theta_{\mu}(y)}},$ (18)
and in the case of the sine window
$A_{\mu}(y)$ $=$ $\frac{\sqrt{2}}{N}\frac{(1+e^{iE_{\mu}TN})e^{i\Theta_{\mu}(y)}\textrm{sin}\frac{\pi}{N}}{(1-e^{i(\Theta_{\mu}(y)+\frac{\pi}{N})})(1-e^{i(\Theta_{\mu}(y)-\frac{\pi}{N})})},$ (19)
where $\Theta_{\mu}(y)=E_{\mu}T-\frac{2\pi}{N}y$. Using the inequalities
$|1-e^{i\theta}|\leq 2$ (20)
and
$\frac{1}{|1-e^{i\theta}|}\leq\frac{\pi}{2|\theta\ \textrm{mod}[-\pi,\pi)|},$ (21)
it can be shown that $P_{\mu}(y)=|A_{\mu}(y)|^{2}$ decays as $O(y^{-2})$ and $O(y^{-4})$ for the rectangular and sine windows, respectively. This fact indicates that when these windows are used as a filter, the error defined in Eq. (14) also decays polynomially in $\max\bigl{\{}|E_{\mu}|^{-1},|\omega_{y_{c}}-E_{\mu}|^{-1}\bigr{\}}$ for $E_{\mu}\leq\omega_{y_{c}}$ or in $\textrm{max}\bigl{\{}|E_{\mu}-\omega_{y_{c}}|^{-1},|\frac{2\pi}{T}-E_{\mu}|^{-1}\bigr{\}}$ for $E_{\mu}>\omega_{y_{c}}$.
The properties of probability distribution for the Kaiser window have been investigated in detail [41, 27, 28]. As can be seen in Fig. 3(c), there is a clear separation between the high-probability region called mainlobe, which is confined in $\approx 2\alpha$ QPE frequency grid points near the peak center, and the rather flat region outside the mainlobe, called sidelobes. The sum of $P_{\mu}(y)$ in the sidelobes gives the amount of the leakage $\epsilon$, which is related to $\alpha$ in the leading order by [41]
$\alpha=O\left(\log\frac{1}{\epsilon}\right).$ (22)
To assess the error in the Kaiser window-based filter, we plot in Fig. 4 the quantity
$\epsilon_{\textrm{max}}=\max_{E_{\mu}}\left[1-\sum_{y}{}^{{}^{\prime}}P_{\mu}(y)\right],$ (23)
where the prime indicates that the sum is over the $\lceil 2\alpha+1\rceil$ highest-probability points near the peak center $E_{\mu}$. This quantity is usually larger than $\Delta\mathcal{R}_{\mu}$ in Eq. (14) and provides an estimate of the maximum error in the Kaiser window-based filter for the states with $\textrm{min}\{E_{\mu},\omega_{y_{c}}-E_{\mu}\}>\omega_{\lceil 2\alpha\rceil}$ or $\textrm{min}\{E_{\mu}-\omega_{y_{c}},\frac{2\pi}{T}-E_{\mu}\}>\omega_{\lceil 2\alpha\rceil}$. The error decreases exponentially in $\alpha$ as expected, and the result has a very weak dependence on the number of frequency grid points $N$.
II.3 QPE-based filters for the three windows
Figure 5: (a) Weight-renormalization factor (filter) $\mathcal{R}_{\mu}$ (Eq. (9)) for the rectangular and the sine windows (top) and for the Kaiser window with $\alpha=3$ and $6$ (bottom). (b) Deviations of $\mathcal{R}_{\mu}$ from the ideal filter, $\Delta\mathcal{R}_{\mu}$ given in Eq. (14), on the logarithmic scale for $E_{\mu}\leq 2$ (left) and $E_{\mu}\geq 5$ (right). The vertical dotted line in the left panel shows the position of $\omega_{y_{c}}$, and the filled squares indicate the positions of the QPE frequency grid points $\omega_{y}$.
Figure 5(a) shows $\mathcal{R}_{\mu}$ (given in Eq. (9)) as a function of $E_{\mu}$ for the three window functions described in the previous section. The parameters used are $T=1$, $n=6$ ($N=64$) and $y_{c}=15$, corresponding to extracting the states in approximately one fourth of the entire energy range. The rectangular-window-based filter exhibits an oscillating behavior known as the Gibbs phenomenon [43], which originates from the discontinuity of the truncated frequency spectrum. In the case of the other two windows this oscillation is suppressed at this scale. As shown in the insets of Figs. 5(a) and 5(b), compared to the rectangular and sine window-based filters, the Kaiser window-based filters have a larger transitional region where $\mathcal{R}_{\mu}$ deviates from one and zero. This reflects the fact that the probability distribution of the Kaiser window has a broader mainlobe as shown in Figs. 3(b) and 3(c).
In Fig. 5(b), the errors of these filters, $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), are plotted together with the QPE frequency grid points (filled squares). The sine window result also shows oscillation, though of smaller magnitude than the rectangular window. By analyzing $P_{\mu}(y)$ [31], it can be shown that these two filters are exact (i.e. $\Delta\mathcal{R}_{\mu}=0$) for $E_{\mu}=\omega_{y^{\prime}}$ and $E_{\mu}=\omega_{y^{\prime}+1/2}$ for the rectangular and sine windows, respectively, with $y^{\prime}$ any integer.
The Kaiser window-based filter has the following properties:
$1-\epsilon$$\leq$ $\mathcal{R}_{\mu},\quad\mbox{for}\quad E_{\mu}\in\left[\frac{\delta}{2},E_{\textrm{targ}}+\frac{\delta}{2}\right],$ (24)
$\mathcal{R}_{\mu}$$\leq$ $\epsilon,\quad\mbox{for}\quad E_{\mu}\in\left[E_{\textrm{targ}}+\frac{3}{2}\delta,\frac{2\pi}{T}-\frac{\delta}{2}\right].$ (25)
Here $E_{\textrm{targ}}<\omega_{y_{c}}$ is the width of the target energy interval, and $\delta$ is the width of the transitional region near $E_{\mu}=0$, $\omega_{y_{c}}$, and $\omega_{N-1}$. The total width of the latter region is $2\delta$. By comparing the Kaiser window results and the QPE frequency grid points, $E_{\textrm{targ}}$ and $\delta$ can be written approximately as
$\delta$$\approx$ $\omega_{\lceil 2\alpha\rceil}=\frac{2\pi}{NT}\lceil 2\alpha\rceil,$ (26)
$E_{\textrm{targ}}$$\approx$ $\omega_{y_{c}}-\delta.$ (27)
From Eq. (26) and using the fact that $\alpha$ and $\epsilon$ are related as $\alpha=O(\log\epsilon^{-1})$, one can obtain
$N\approx\frac{2\pi}{\delta T}2\alpha=O(\delta^{-1}\log\epsilon^{-1}).$ (28)
This is the number of time evolution operations $U_{HT}$, required to filter the state with given error $\epsilon$.
II.4 The Gibbs phenomenon and the $\sigma$-factor
As seen in the previous section, the QPE-based filter $\mathcal{R}_{\mu}$ generally exhibits oscillation known as the Gibbs phenomenon [43]. To analyze this effect, let us first rewrite $\mathcal{R}_{\mu}$ as
$\mathcal{R}_{\mu}=\sum_{j=-(N-1)}^{N-1}\tilde{g}_{y_{c}}(j)\sigma_{j}e^{-iE_{\mu}Tj},$ (29)
where
$\tilde{g}_{y_{c}}(j)=\frac{e^{i\frac{\pi}{N}y_{c}j}}{N}\frac{\textrm{sin}\frac{\pi(y_{c}+1)j}{N}}{\textrm{sin}\frac{\pi j}{N}}$ (30)
is the discrete Fourier transform of the discretized step function $g_{y_{c}}(y)$:
$g_{y_{c}}(y)$ $=$ $\Bigl{\{}\begin{array}[]{ll}1&\quad\mbox{for}\quad 0\leq y\leq y_{c}\\ 0&\quad\mbox{for}\quad y_{c}<y\leq N-1\end{array}\Bigr{.}.$ (33)
For $N\gg j$, Eq. (30) becomes $\tilde{g}_{y_{c}}(j)\to e^{i\frac{z_{c}}{2}j}\frac{\textrm{sin}\frac{z_{c}}{2}j}{\pi j}$ with $z_{c}=\frac{2\pi}{N}y_{c}$, which is the Fourier series expansion coefficient of the following $2\pi$-periodic function $G_{z_{c}}(z)$:
$G_{z_{c}}(z)$ $=$ $\Bigl{\{}\begin{array}[]{ll}1&\quad\mbox{for}\quad 0\leq z\leq z_{c}\\ 0&\quad\mbox{for}\quad z_{c}<z\leq 2\pi\end{array}\Bigr{.}$ (36)
Equation (29) has the same form as the truncated Fourier series expansion of the step function with a factor $\sigma_{j}$, which is constructed from the QPE window function $a_{j}$
$\sigma_{j}=\sum_{j^{\prime}=0}^{N-1-|j|}a_{|j|+j^{\prime}}a_{j^{\prime}}.$ (37)
In the theory of the Gibbs phenomenon [43], this $\sigma$-factor is called a “filter” (not to be confused with filtering of states discussed in earlier sections) and serves as a smoothing factor that improves the convergence of the truncated Fourier series. Several filters have been proposed in the classical Fourier spectral analysis. Some examples are the Lanczos filter and the raised cosine filter [43, 44].
The explicit form of $\sigma_{j}$ for the rectangular window is
$\sigma_{j}=\frac{N-|j|}{N}$ (38)
and for the sine window
$\sigma_{j}=\frac{1}{N\sin\frac{\pi}{N}}\Bigl{[}\textrm{sin}\frac{\pi|j|}{N}\cos\frac{\pi}{N}+(N-|j|)\cos\frac{\pi|j|}{N}\sin\frac{\pi}{N}\Bigr{]}.$ (39)
In the limit $N\to\infty$, Eqs. (38) and (39) become
$\sigma(x)=1-|x|$ (40)
and
$\sigma(x)=\frac{1}{\pi}\sin(\pi|x|)+(1-|x|)\cos(\pi x),$ (41)
respectively, with $x=j/N$.
In Fig. 6, we plot the above two forms as a function of $x=j/N$ as well as the Kaiser window result calculated numerically. In the rectangular window case, $\sigma_{j}$ has a triangular structure with a singularity at $j=0$. In the case of the sine and the Kaiser windows, $\sigma_{j}$ is more localized at $x=0$, therefore it filters out the Fourier expansion coefficients $\tilde{g}_{y_{c}}(j)$ in Eq. (29) for large $j$. This accelerates the convergence of the Fourier series expansion, and yields a smoother $\mathcal{R}_{\mu}$, as observed in Fig. 5. More detailed analysis of the $\sigma$-factor may offer a way to improve the window function, but is beyond the scope of the present work.
Figure 6: Comparison of the $\sigma$-factors defined in Eq. (37) as a function of $x=j/N$ for $N=64$.
II.5 Comparison with QETU
Figure 7: Quantum circuit for QETU. Here $U_{HT/2}=e^{iHT/2}$.
In this section, we compare the QPE-based and the QETU-based filters. Figure 7 shows the quantum circuit for QETU [12] with one ancilla qubit. It consists of $d+1$ rotation gates with symmetric $d+1$ phase angles $(\varphi_{0},\varphi_{1},\dots,\varphi_{1},\varphi_{0})$ and $d$ blocks of $U^{\dagger}_{HT/2}=e^{-iHT/2}$ and $U_{HT/2}$ (note that in our case $HT$ is defined in the interval $[0,2\pi]$). The output of this circuit after measuring the ancilla qubit is $\frac{f_{d}(w)|\psi\rangle}{||f_{d}(w)|\psi\rangle||}$, where $f_{d}(w=\cos\frac{HT}{4})$ is an even polynomial of degree $d$. It has been shown in Ref. [12] that the QETU-based filter can be constructed with $d=O(\delta^{-1}\log\epsilon^{-1})$ operations of $U^{\dagger}_{HT/2}$ and $U_{HT/2}$. This is the same order as the Kaiser window-based QPE filter, although in the case of the Kaiser window the extra cost of the window function preparation needs to be taken into account. For example, in the approach of Ref. [42], a system-independent cost of $O(n\sqrt[4]{\alpha+1}(\alpha+\log\epsilon^{-1}))=O(n(\log\epsilon^{-1})^{5/4})$ has to be added.
Our interest here is to compare the performance of the two filtering approaches in terms of the number of queries to $U_{HT}$, which has the most significant impact on the circuit depth. For this purpose, we perform a numerical simulation to estimate the number of queries in both approaches for a given set of parameters $E_{\textrm{targ}},\delta$, and $\epsilon$. Although QETU can produce a polynomial function satisfying Eqs. (24) and (25), here we consider a more simple and practical filtering function $F(E)$, which corresponds to $\mathcal{R}_{\mu}$ in Eq. (9), that satisfies the following conditions:
$|F(E)|$$\leq$ $1\quad E\in[0,\frac{2\pi}{T}]$ (42)
$1-F(E)$$\leq$ $\epsilon\quad E\in[0,E_{\textrm{targ}}]$ (43)
$|F(E)|$$\leq$ $\epsilon\quad E\in[E_{\textrm{targ}}+2\delta,\frac{2\pi}{T}].$ (44)
This function has the same target region $E_{\textrm{targ}}$ and the transitional region $2\delta$ as in Eqs. (24) and (25).
Note that the filtering function $F(E)$ is applied to the squared amplitude for each state as $|C_{\mu}|^{2}\to F(E_{\mu})|C_{\mu}|^{2}$, rather than $C_{\mu}$ itself. We therefore construct the function $f(w)$ such that
$|f(w)|$$\leq$ $1\quad w\in[0,1]$ (45)
$f(w)$ $=$ $c\quad w\in[w_{+},1]$ (46)
$f(w)$ $=$ $0\quad w\in[0,w_{-}]$ (47)
where
$c$ $=$ $\frac{1}{2}\Bigl{[}1+\sqrt{1-\epsilon}\Bigr{]},$ (48)
$w_{+}$ $=$ $\cos\frac{E_{\textrm{targ}}}{4}T,$ (49)
$w_{-}$ $=$ $\cos\frac{(E_{\textrm{targ}}+2\delta)}{4}T.$ (50)
This function is related to $F(E)$ satisfying the conditions Eqs. (42-44) as $F(E)=|f(\cos\frac{ET}{4})|^{2}$. This function is approximated by Chebyshev polynomials up to even degree $d$ as
$f(w)\approx f_{d}(w)=\sum_{l=0}^{d/2}c_{2l}T_{2l}(w),$ (51)
where $T_{2l}(w)$ and $c_{2l}$ are the $2l$-th Chebyshev polynomials and expansion coefficients, respectively. We numerically determine the smallest $d$ in which $F_{d}(E)=|f_{d}(\textrm{cos}\frac{ET}{4})|^{2}$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$. The expansion coefficients $c_{2l}$ are calculated with the convex-optimization-based approach proposed in Ref. [12] and using the qsppack library [45, 46, 47]. We use $M$ grid points $w_{m=0,1,\dots,M-1}$ in $[0,1]$ and solve the following optimization problem
$\min_{\{c_{2l}\}}\max\Bigl{[}$ $\max_{w_{m}\in[w_{+},1]}|f_{d}(w_{m})-c|,$
$\max_{w_{m}\in[0,w_{-}]}|f_{d}(w_{m})|\Bigr{]}$
s.t. $|f_{d}(w_{m})|\leq 1\quad(m=0,1,\dots,M-1).$ (52)
After the optimal coefficients $c_{2l}$ are obtained with $M$ sampled grid points, we check whether the function $F_{d}(E)$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$ by combining a grid search with a fine grid of $100,000$ points and an optimization routine using the scipy.optimization module [48]. As an example of the calculated results, $f_{d}(w)$ and $F_{d}(E)$ obtained with degree $144$ and $\epsilon=0.01$ are shown in Figs. 8(a) and 8(b).
Figure 8: Polynomial approximations of (a) $f(w)$ and (b) $F(E)=|f(\cos\frac{ET}{4})|^{2}$ with degree $144$ and $\epsilon=0.01$. The parameters used are $E_{\textrm{targ}}=1.4430,2\delta=0.2556$, and $T=1$. The two vertical dotted lines indicate the positions of $w_{\pm}$ in (a) and $E_{\textrm{targ}}$ and $E_{\textrm{targ}}+2\delta$ in (b). The inset in (b) shows a magnified view.
Figure 9: Minimum number of queries to $U_{HT}$ in QETU obtained with $M=800$, $3200$, and $12800$ for three sets of ($E_{\textrm{targ}},\delta$). The corresponding values in the QPE-based filters are also shown for comparison. The straight lines show the results of the linear regression. (a) $E_{\textrm{targ}}=1.0589$ and $1.02322\delta=$ 1.0232. (b) $E_{\textrm{targ}}=1.4430$ and $0.25562\delta=$ 0.2556. (c) $E_{\textrm{targ}}=1.5069$ and $0.12792\delta=$ 0.1279. In (c) the results for $\epsilon=10^{-4}$ and $\epsilon=10^{-5}$ with $M=800$ are not calculated as there are less sampled points than the required degree $d$.
Figure 9 shows the calculated number of queries to $U_{HT}$, which is half of the minimum degree $d$, for three different sets of $E_{\textrm{targ}},\delta$. Here $E_{\textrm{targ}}$ and $\delta$ are obtained numerically from the results of the Kaiser window-based filter (Fig. 5 (b)) with $\alpha=3$ and $n=6,8,9$, using $y_{c}=2^{n-2}-1$ and $\epsilon=10^{-7}$ (see also Fig. 4). The number of sampled points $M$ is a hyperparameter, and we compare three cases, $M=800,3200$ and $12800$. In the figures, $\epsilon$ is varied from $10^{-2}$ to $10^{-5}$. For $\epsilon<10^{-5}$ we have difficulty obtaining reliable results. The errors in the QPE-based filters with the three window functions are also shown for comparison.
In QETU the number of queries increases linearly with $\log(\epsilon^{-1})$ as expected, and decreases as the number of sampled points $M$ increases. The latter trend can be understood by noting that with a small $M$ the optimized coefficients tend to overfit to the sampled data points. Interestingly, for the three cases studied, the QETU results extrapolated to $\epsilon=10^{-7}$ are very close to the results of the Kaiser window-based filter (crosses). This can potentially be understood by noting that the Kaiser window is an approximation of the prolate spheroidal window, which is designed to maximize the probability of the mainlobe of the probability distribution $P_{\mu}(y)$ [26]. This construction is approximately equivalent to minimizing the error of the QPE-based filter $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), potentially explaining the similarities between the QETU-based filter obtained from the optimization (Eq. (52)) and the Kaiser window-based filter. Although the extra cost of embedding the window function $a_{j}$ and $O(\log\bigl{[}\delta^{-1}\log\epsilon^{-1}\bigr{]})$ ancilla qubits may prevent the Kaiser window-based filter from being the method of choice, our result yields insights into the nature of the Kaiser window and its potential usefulness as a QPE-based alternative to QETU and other QSP-based filtering approaches.
In relation to this study, we note that the Chebyshev expansion coefficients $c_{2l}$ are sometimes calculated via the procedure proposed by Low and Chuang [49]. Their procedure approximates the discontinuous sign function by the error function, which is in turn approximated by using the known polynomial expansion of the exponential decay function. The corresponding expansion coefficients are obtained by evaluating the modified Bessel functions of the first kind [49]. We also estimate the number of degrees in this procedure using the analytic formula derived from it [50, 51], and find that it generally requires larger number of queries compared to the current convex-optimization-based results. This is understandable, as the convex-optimization-based approach does not assume any specific form for the polynomial function, therefore the expansion coefficients ($c_{l}$) are subject to fewer restrictions. However, getting the coefficients via numerical optimization becomes more difficult as one requires higher accuracy, therefore for very small $\delta$ or $\epsilon$ one has to rely on non-optimization-based approaches such as Low and Chuang’s procedure. In such cases, the QPE-based filter with the Kaiser window may be a competitive alternative.
III Application to calculations of low-energy spectra
Figure 10: Quantum circuit for the QPE-based spectral simulation with filtering.
III.1 QPE-based spectral simulation
Finally, we consider an application of the present approach to the QPE-based calculation of the dynamical response function [34, 29, 30, 31, 32, 33, 35]. The objective is to calculate
$S_{B}(\omega)=\sum_{\mu}|\langle\phi_{\mu}|B|\phi_{0}\rangle|^{2}\delta(\omega-\Delta E_{\mu}),$ (53)
where $B$ is an operator which is not necessarily unitary, and $\Delta E_{\mu}=E_{\mu}-E_{0}$. The quantity $S_{B}(\omega)$ is related to the Fourier transform of the dynamical response function or the correlation function in the time domain and can be compared with various experimental spectra.
In the QPE-based approach, $S_{B}(\omega)$ is obtained using the quantum circuit in Fig. 1 by modifying the state preparation operation $U_{\psi}$ to block-encode $B|\phi_{0}\rangle$ with $a$ ancilla qubits as
$\Bigl{[}\langle 0|^{\otimes a}\otimes I\Bigr{]}U_{\psi}|0\rangle^{\otimes a}|0\rangle^{\otimes s}$ $=$ $B|\phi_{0}\rangle$ (54)
$=$ $\sum_{\mu}\langle\phi_{\mu}|B|\phi_{0}\rangle|\phi_{\mu}\rangle,$
and by replacing $H$ in $U_{HT}$ with $H-E_{0}$ if necessary. By collecting the measurement outputs in Fig. 1, the following approximation of $S_{B}(\omega)$ at discrete frequency points $\omega_{y}$ is obtained:
$\tilde{S}_{B}(\omega_{y})=\sum_{\mu}|\langle\phi_{\mu}|B|\phi_{0}\rangle|^{2}P_{\mu}(y).$ (55)
Figure 10 shows the quantum circuit with filtering applied prior to the QPE operation. Such filtering is useful in cases where one is interested in some specific small energy region, as in the case of virtual screening in materials science. The first step of this circuit performs filtering with $n$ ancilla qubits followed by QAA, as indicated by the gate labeled “Filter + QAA” in Fig. 10. Here we apply the proposed QPE-based filtering. The ancilla register can either be traced out or reused after measurement. The second step is to perform another QPE to obtain $\tilde{S}_{B}(\omega_{y})$ with $n^{\prime}$ ancilla qubits and time step $T^{\prime}$. The parameters $n^{\prime}$ and $T^{\prime}$ as well as the window function in $U_{\textrm{in}}^{\prime}$ used in the second step can be chosen independently to the ones used in the filtering step.
The final frequency resolution of this approach is $\Delta\omega^{\prime}=\frac{2\pi}{N^{\prime}T^{\prime}}$, with $N^{\prime}=2^{n^{\prime}}$. The parameters $n$ and $T$ used in the first filtering step can be smaller than those used in the second step, as in the coarse QPE approach proposed in Ref. [15]. In a naive application of QAA without filtering, the total number of queries to $U_{HT}$ is $O(p_{0}^{-1/2}\beta N^{\prime})$, where $\beta=T^{\prime}/T$ and $p_{0}$ is the probability of measuring $\tilde{S}_{B}(\omega_{y})$ in the energy range of interest. With filtering, the total number of queries is reduced to $O(p_{0}^{-1/2}N+\beta N^{\prime})$, with $N=O(\delta^{-1}\log\epsilon^{-1})$ for the case of Kaiser window-based filter. This two-step approach therefore allows for the efficient calculation of the energy spectrum of the specified small energy region when $\beta N^{\prime}\gg N$ and can be useful to get fine structures of the spectrum at high resolution.
III.2 Density of states of MnO
Figure 11: Occupied part of the DOS of MnO (a) for the entire energy region and (b, c) near the cutoff after filtering using the rectangular window (b) and the Kaiser window with $\alpha=3$ (c). The solid curves show the DOS calculated with a fine QPE frequency grid, and the vertical dotted lines indicate the position of the cutoff frequency $\omega_{y_{c}}$. The states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$ are filtered out. The filled squares in (b) and (c) indicate the QPE frequency points used in filtering. Note that the Fermi level is not set to zero in this calculation.
Figure 12: (a) Error of the calculated DOS after filtering. (b) Contribution from the states with $\varepsilon_{\mathbf{k}\nu}\geq\omega_{y_{c}}$. (c) Contribution from the states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$. The vertical dotted lines indicate the position of $\omega_{y_{c}}$, and the filled squares in (b) and (c) indicate the QPE frequency points used in filtering.
To see the performance of the proposed QPE-based filtering for realistic spectra, we apply it to the calculation of the DOS [36] of the type-II antiferromagnetic MnO. This material is a prototype of transition metal monoxides, where strong electron correlation between the $3d$ orbitals of the transition metal atoms in a narrow energy region plays a crucial role in determining their electronic and magnetic properties. The insulating nature of the type-II antiferromagnetic MnO is explained by the combination of the magnetic ordering of Mn and the coupling between Mn $3d$ and O $2p$ orbitals [52, 53], and a number of band-structure calculations have been performed to understand the exact electronic structure of this system and also as a benchmark of a newly proposed algorithm  [54, 55, 56, 57, 58, 59, 60].
In this work we focus on the occupied part of the DOS of this material, $D(\omega)$, within the one-particle approximation:
$D(\omega)=\sum_{\mathbf{k}\nu}^{\textrm{occupied}}\delta(\omega-\varepsilon_{\mathbf{k}\nu}),$ (56)
where $\varepsilon_{\mathbf{k}\nu}$ are one-particle orbital energies, and $\mathbf{k}$ and $\nu$ are the Bloch wavevectors and orbital (or band) indices [36], respectively. Although this DOS itself can be calculated efficiently on a classical computer, we use it as an approximation of the spectral function for interacting many-electron systems, which is hard to calculate classically [30, 32, 31]. The DOS and spectral function are related to the experimental spectra obtained in photoemission spectroscopy.
In this work we prepare $\varepsilon_{\mathbf{k}\nu}=E_{\mu=(\mathbf{k}\nu)}$ classically using density functional theory [61] and set
$B=\sum_{\mathbf{k}\nu}\hat{c}_{\mathbf{k}\nu}^{\dagger}\hat{c}_{\mathbf{k}\nu},$ (57)
where $\hat{c}_{\mathbf{k}\nu}^{\dagger}$ and $\hat{c}_{\mathbf{k}\nu}$ are the electron creation and annihilation operators, respectively. The DOS is calculated via Eq. (55) using a noiseless simulator. The input data is prepared using QuantumEspresso [62, 63] with the pseudopotentials [36] from its pseudopotential database [64]. The rocksalt structure is assumed with lattice constant $a=4.43$. The generalized gradient approximation (GGA) [65] is used for the exchange and correlation functional, and the kinetic energy cutoff for wavefunctions used is $47$ Ry. The one-particle energies $\varepsilon_{\mathbf{k}\nu}$ are calculated with $32\times 32\times 32$ $\mathbf{k}$-points after obtaining the electron charge density self-consistently. Although the number of states in this calculation is prohibitively large for early fault-tolerant quantum computers, the resulting DOS yields a continuous spectrum that helps us to evaluate the efficacy of the proposed approach.
Figure 11(a) shows the DOS of this system without filtering, which is obtained with the sine window function for $U_{\textrm{in}}^{\prime}$. The sharp peaks below $\approx-0.2$ hartree from the Mn $3s$ and $3p$ orbitals and O $2s$ orbital play a minor role in determining the properties of this material. The broader spectra above $\approx+0.2$ hartree consist mainly of O $2p$ and Mn $3d$ orbitals. Our interest here is to obtain the DOS of the Mn $3d$ orbitals at high resolution. We therefore set a cutoff $\omega_{y_{c}}\approx 0.378$ as shown in the vertical dotted line in Fig. 11(a), and filter out the states below this threshold. In the calculation, we change the sign of $\varepsilon_{\mathbf{k}\nu}$ and apply a low-pass filter without modifying the current formalism.
The spectrum above the cutoff consists mainly of Mn $3d$ states. As can be seen in Fig. 11(b), due to the crystal field effect it splits into two main peaks [53, 52]: the lower-lying $t_{2g}$ peak and the higher-lying $e_{g}$ peak. The $e_{g}$ peak lies approximately from $0.42$ to $0.46$ hartree, and we set this energy region as the target energy range. We add a padding of 0.5 hartree above the $e_{g}$ states. Figures 11(b) and 11(c) show the DOS (obtained via Eq. (55)) after QPE-based filtering in open circles, calculated with the rectangular window (Fig. 11(b)) and with the Kaiser window using $\alpha=3$ (Fig. 11(c)). Note in both cases the sine window function is used in the second step. The parameters used are $n=n^{\prime}=6$ ($N=N^{\prime}=128$) and $T=2,T^{\prime}=20$, and the coarse QPE frequency grid points used in the filtering step are shown by the filled squares.
At this scale both the rectangular and the Kaiser windows efficiently filter out the states below the cutoff, which mainly consist of O $2p$ states, and both results yield higher-resolution spectra of the target ($e_{g}$) as well as $t_{2g}$ states, in good agreement with the original DOS. The Kaiser window result shows a better agreement with the original DOS without filtering, as expected.
Figure 12(a) shows the error of the calculated DOS after filtering for the three window functions. This error is obtained as the deviation from the unfiltered high-resolution DOS calculated with the states above $\omega_{y_{c}}$,
$\tilde{D}(\omega)=\sum_{\mathbf{k}\nu}^{\textrm{occupied},\varepsilon_{\mathbf{k}\nu}\geq\omega_{y_{c}}}P_{\mu=(\mathbf{k}\nu)}(y).$ (58)
The error is largest in the rectangular window result, and the error in the Kaiser window result decreases after ±$\pm\alpha$ frequency grid points used in filtering from the cutoff frequency $\omega_{y_{c}}$, in accordance with the observation in the previous section. Switching from the rectangular window to the Kaiser window reduces the error in the target energy region about $0.42$ to $0.46$ hartree, which corresponds to the $e_{g}$ orbital energies (see Figs. 11(b) and 11(c)), by about six orders of magnitude.
In Figs. 12(b) and 12(c) we decompose the error into the contribution from the states with (b) $\varepsilon_{\textbf{k}\nu}\geq\omega_{y_{c}}$ (i.e., Mn $t_{2g}$ and $e_{g}$) and (c) $\varepsilon_{\textbf{k}\nu}<\omega_{y_{c}}$. By comparing Figs. 12(a) and 12(b), it can be seen that the dominant error in the target region ($e_{g}$ states of Mn) originates from the states with $\varepsilon_{\textbf{k}\nu}\geq\omega_{y_{c}}$. Similarly, Fig. 12(c) shows the states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$ are the main source of the error in region $\omega<\omega_{y_{c}}$. In Fig. 12(c) some extra low-intensity peaks are seen in the region $\omega\geq\omega_{y_{c}}$. These peaks originate from the low-lying states (see Fig. 11(a)) whose energies $\varepsilon_{\mathbf{k}\nu}$ are related to $\omega$ via $\omega=\varepsilon_{\mathbf{k}\nu}+\frac{2\pi}{T^{\prime}}r$ with $r$ a positive integer, as phases in QPE are defined modulo $2\pi$. This error is also mostly suppressed in the Kaiser window result.
In summary, this calculation demonstrates that the proposed two-step approach provides an efficient way of calculating high-resolution spectra of the specified small energy region. This approach is applicable to general discrete and continuous spectral simulations of quantum many-body systems, which are hard to calculate via classical means. The Kaiser window-based filter provides an accurate and controllable approach, with tunable parameters $N$ and $\alpha$ based on the specified target energy interval and the tolerance error $\epsilon$ using Eqs. (26) and (27).
IV conclusions
In this work, we have investigated the properties of the QPE-based filtering approach for three different window functions. We have shown that the results with the conventional rectangular window exhibit an oscillation known as the Gibbs oscillation, and this oscillation is suppressed when using the sine and the Kaiser windows. We have also shown that this difference is explained by the $\sigma$-factor (smoothing factor) appearing in the theory of the Gibbs oscillation, the tuning of which may contribute to the future improvement of window functions for QPE.
We have found that, under the same conditions the number of queries to $U_{HT}$ required in the Kaiser window-based filter is comparable to that in QETU after optimization of phase angles. This indicates that the form of the Kaiser window function is nearly optimal for filtering and can be used as a potential alternative to QETU, although one should keep in mind that there is an additional cost associated with window state preparation.
As an application of the current approach, we have considered the combination of the filtering technique to the QPE-based spectral calculation. We have also performed a numerical simulation of the DOS of type-II antiferromagnetic MnO as an approximation of the spectral function. Our results have shown that the QPE-based filtering efficiently extracts the energy range of interest, in this case the $e_{g}$ orbitals of Mn. This calculation has also shown that even the simplest rectangular window sufficiently reproduces the peak structures of this system. These results indicate that, combined with efficient ground-state preparation and time evolution techniques, the proposed algorithm can be used as a simple QPE-based scheme complementary to QSP-based approaches for future materials simulations on a fault-tolerant quantum computer.
Acknowledgements.
We thank Hiroshi Yano, Jumpei Kato, Kohei Oshio, and Pawan Poudel for useful discussions. This work was supported by MEXT Quantum Leap Flagship Program Grants No. JPMXS0118067285 and No. JPMXS0120319794. K.W. was supported by JSPS KAKENHI Grant Number JP24KJ1963. K.S. acknowledges support from Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from JSPS, Japan.
Appendix A The case of classical post-selection
Here we provide a detailed explanation of the classical filter based on post-selection. Recall that the output state of the QPE circuit (Fig. 1) before the measurement is given by
$|\Psi\rangle$ $=$ $\sum_{y=0}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle$
$=$ $\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}|\chi_{\mu}\rangle|\phi_{\mu}\rangle+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle.$
When the measurement outcome is $y$, the state after the measurement becomes
$\frac{\sum_{\mu}C_{\mu}A_{\mu}(y)|{\phi_{\mu}}\rangle}{\sqrt{\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y)|^{2}}},$ (59)
and the squared amplitude of state $|\phi_{\mu}\rangle$ is
$\frac{|C_{\mu}|^{2}|A_{\mu}(y)|^{2}}{\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y)|^{2}}.$ (60)
The expectation value of this squared amplitude is obtained by multiplying Eq. (60) by the conditional probability
$\frac{\sum_{\mu}|C_{\mu}|^{2}|A_{\mu}(y)|^{2}}{\sum_{y^{\prime}=0}^{y_{c}}\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y^{\prime})|^{2}}$ (61)
and summing over $y=0,1,\dots,y_{c}$. This cancels out the denominator of Eq. (60) and yields $|C_{\mu}|^{2}\mathcal{R}_{\mu}$ up to a constant factor. Therefore Eq. (11) holds for the classical case as well as the quantum (coherent) case.