### v3

# Quantum phase estimation based filtering: performance analysis and application to low-energy spectral calculation

Abstract
Filtering refers to the process of using algorithms to enhance desired or mask undesired states in quantum computing, and is essential for preparing or isolating quantum many-body states. One class of filters is based on quantum phase estimation (QPE). Such QPE-based filters have a simple circuit implementation, but their performance remains unclear, particularly when compared to other filtering methods such as those with quantum signal processing (QSP). In this work, we investigate the performance of QPE-based filters with three commonly used window functions: rectangular window, sine window, and Kaiser window. We analyze how choice of window affects results; one noticeable observation is that the oscillation of the filter (the Gibbs phenomenon) seen in the conventional rectangular window result can be mitigated when the sine and Kaiser windows are applied. Furthermore, we compare the performance of QPE-based methods to quantum eigenvalue transformation of unitary matrices with real polynomials (QETU), and find that the Kaiser window-based filter and QETU result in a similar number of queries to the Hamiltonian time evolution operation. QPE-based filtering methods may therefore be an alternative to signal-processing-based methods. Finally, as an application of the QPE-based filter, we propose a two-step QPE algorithm for low-energy spectral simulations, composed of a coarse grid for filtering and a fine grid for obtaining final high-resolution spectra. As a benchmark of the proposed scheme for realistic continuous spectra, we present the density-of-states (DOS) calculation of antiferromagnetic type-II MnO in a one-particle approximation.
I introduction
Calculation of electronic structure of complex materials is one of the most promising industrial applications of quantum computing. The last decade has seen rapid development of quantum-classical hybrid approaches that leverage currently available noisy intermediate-scale quantum computers, such as the variational quantum eigensolver [Peruzzo2014] and sampling-based approaches [kanno2023quantumselectedconfigurationinteractionclassical, robledomoreno2024chemistryexactsolutionsquantumcentric]. As advances in fault-tolerant hardware and error correction techniques continue, fully-quantum algorithms [zhang2025faulttolerantquantumalgorithmsquantum] such as quantum phase estimation (QPE) [kitaev1995quantummeasurementsabelianstabilizer] promise true quatum advantage in the form of exponential speed up over classical computers.
However, even with fault-tolerant hardware, preparing the precise ground state and resolving low-energy spectra are crucial, as many phenomena of industrial and academic importance occur at the small energy scales between the electronic ground state and low-lying excited states. For example, semiconductor band gaps are typically on the order of $1$ eV [yu2010fundamentals], and the energy scale of many relevant types of chemical reactions, high-temperature superconductivity [brooks2007handbook], and spintronics [RevModPhys.76.323, spintronics2] are in the meV range. On the other hand, the total energy and the L2 norm of the Hamiltonian $\|H\|_{2}$ are extensive, and even for small molecules exceed 1010010-100 eV. This difference in energy scales not only makes the ground-state preparation difficult, as shown in recent sample-based calculations [robledomoreno2024chemistryexactsolutionsquantumcentric, D5CP02202A, yu2025quantumcentricalgorithmsamplebasedkrylov], but also makes it hard to resolve small low-energy excitations in quantum algorithms such as QPE.
To overcome this fundamental problem, quantum filtering algorithms have been developed to obtain the desired quantum state from a given input state. Examples of the ground-state filtering include the cosine [10.1063/1.5027484], Gaussian [PhysRevA.106.032420, Wang2022statepreparation], and Lorentzian [Irmejs2024efficientquantum]-based filtering schemes. There are also filtering methods based on quantum signal processing (QSP) [Low2019hamiltonian, PhysRevLett.118.010501], such as quantum singular value transformation (QSVT) [10.1145/3313276.3316366, PRXQuantum.2.040203, Lin2020nearoptimalground] and quantum eigenvalue transformation of unitary matrices with real polynomials (QETU) [PRXQuantum.3.040305]. However, in all QSP-based methods one has to find a set of single qubit phase angles to construct the target function on a classical computer. This procedure requires two steps. First, one needs to find a good polynomial approximation of the target function. Second, the phase factors realizing this polynomial function must be determined numerically. Although classical algorithms have been developed for the angle-finding procedure [Haah2019product, chao2020findinganglesquantumsignal, PhysRevA.103.042419, Wang2022energylandscapeof, Ying2022stablefactorization], it is worth looking for a simpler alternative filtering approach that avoids this procedure.
Here we choose to study filters based on QPE, a class of algorithms that leverages the fact that eigenenergies of the system are approximately encoded in the ancilla register in binary expansion form. In an early work, Poulin and Wocjan [PhysRevLett.102.130503] detail a ground-state preparation scheme based on a repeated backward operation of QPE, and in Ref. [Jensen_2021], P. W. K. Jense et al. propose a general systematic approach of amplifying quantum states within a specified energy interval. In addition, a recent study reported [PRXQuantum.5.040339] a simple model calculation in which their QPE-based filter with a coarse frequency grid was more cost-effective than QETU, suggesting QPE has the potential to outperform QSP-based methods. However, without a thorough understanding of the behavior and properties of QPE-based filters, it is difficult to assess any potential advantages.
We therefore aim to investigate the characteristics of the QPE-based filter through theoretical and numerical analyses, particularly the effect of the initial state of the ancilla register. This initial state, called the window function, is known to affect the performance of QPE [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015, patel2024optimalcoherentquantumphase, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], prompting questions about its role in QPE-based filters. We focus on filters that increase the probability of the states in a specified small energy region without altering their relative amplitudes, rather than increasing the probability of obtaining the ground state of the system as done in many ground-state preparation techniques. This style of filtering has various applications, such as spectroscopy simulations.
We study three commonly used window functions: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. The rectangular window is the most common and used in the conventional QPE algorithm, although it is known to suffer from the spectral leakage problem [1455106, 9762511]. The sine window is known to mitigate the problem and has been applied in quantum metrology [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015] to achieve the sought-after Heisenberg-limit scaling. The Kaiser window [1163349] is proposed as an approximation of the prolate spheroidal window or the Slepian window [6773659, https://doi.org/10.1002/sapm196544199, Imai_2009, patel2024optimalcoherentquantumphase], and recently advantages of using this window over QSVT-based [PRXQuantum.2.040203, greenaway2024casestudyqsvtassessment] and single-ancilla [PRXQuantum.6.020327] phase estimation algorithms have been reported.
First, we analyze the general properties of QPE-based filters. We show that the rectangular window-based filter exhibits an oscillation known as the Gibbs phenomenon, and the oscillation is mitigated when the sine and Kaiser windows are used. We also show that the effect of the window function on the filter can be incorporated into a multiplicative factor. The error in the Kaiser window-based filter decreases exponentially with the QPE frequency grid points, leading to the $O(\delta^{-1}\log\epsilon^{-1})$ query complexity. Here $\delta$ and $\epsilon$ are the width of the transitional region and the error parameter, respectively.
Next, we compare the performance of QPE-based filters and a single-ancilla QSP-based approach QETU [PRXQuantum.3.040305]. The QETU method is known to achieve nearly optimal query complexities for the ground-state energy estimation and the ground-state preparation [PRXQuantum.3.040305], and recently a demonstration on trapped-ion quantum hardware has been reported [karacan2025filterenhancedadiabaticquantumcomputing]. We find that the Kaiser window-based filter and QETU with phase angles determined via convex optimization require a comparable number of Hamiltonian evolution operations for a given set of ($\delta,\epsilon$). This result indicates that the Kaiser window-based filter can be used as a potential alternative to the QETU-based filter. We discuss possible cases where the QPE-based filter can be advantageous compared to QETU.
Finally, we consider an application of the proposed filtering approach to the QPE-based calculation of spectral functions [PhysRevC.100.034610, PhysRevA.101.012330, PhysRevA.110.022618, PhysRevA.110.L060401, fomichev2024simulatingxrayabsorptionspectroscopy, PhysRevA.92.062318, nishi2025demonstrationlogicalquantumphase]; specifically, a two-step QPE approach similar to the one proposed in Ref. [PRXQuantum.5.040339], that uses a coarse QPE grid for filtering and a fine grid for obtaining high-resolution spectra. We evaluate the performance of the proposed scheme by calculating the density-of-states (DOS) of MnO in a one-particle approximation [martin2020electronic] as an example of realistic spectra.
II theory
II.1 General formalism
Figure 1 shows the general form of the quantum circuit for QPE, which consists of $n$ and $s$ qubits for the ancilla and system registers, respectively.
Figure 1: Quantum circuit for QPE.
Here $U_{\textrm{in}}$ constructs the input state for the ancilla register
$U_{\textrm{in}}|0\rangle^{\otimes n}=\sum_{j=0}^{N-1}a_{j}|j\rangle$ (1)
with $N=2^{n}$. The real coefficients $a_{j}$ are related to the window function in classical signal processing, and specific forms of $a_{j}$ are discussed in the next session.
The state-preparation operation $U_{\psi}$ acts on the system register as
$U_{\psi}|0\rangle^{\otimes s}$ $=$ $|\psi\rangle$ (2)
$=$ $\sum_{\mu}C_{\mu}|\phi_{\mu}\rangle,$
where $\phi_{\mu}$ are the eigenstates of the Hamiltonian $H$ with eigenenergies $E_{\mu}$, and $C_{\mu}$ are expansion coefficients normalized as $\sum_{\mu}|C_{\mu}|^{2}=1$. After the state-preparation step, the controlled time evolution $U_{HT}=\textrm{exp}[iHT]$ is performed, where $T$ is a time duration chosen to scale the energy range of the system in $0,2\pi]$. Throughout this work $\hbar=1$. Finally, the inverse of quantum Fourier transform (QFT) is performed. The quantum state before the measurement in Fig. 1 can be represented as
$|\Psi\rangle$ $=$ $\sum_{y=0}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle,$ (3)
where
$A_{\mu}(y)=\frac{1}{\sqrt{N}}\sum_{j=0}^{N-1}a_{j}\textrm{exp}[i(E_{\mu}T-\frac{2\pi}{N}y)j].$ (4)
The probability of measuring $y=0,1,\dots,N-1$ for each state $|\phi_{\mu}\rangle$ is
$P_{\mu}(y)=|A_{\mu}(y)|^{2},$ (5)
and the total probability of measuring $y$ is given as the sum over $\mu$,
$P(y)=\sum_{\mu}|C_{\mu}|^{2}P_{\mu}(y).$ (6)
For $N\to\infty$, each $P_{\mu}(y)$ in Eq. (5) has a peak at $E_{\mu}$ when seen as a function of frequency points
$\omega_{y}=\frac{2\pi}{NT}y.$ (7)
However, since $P_{\mu}(y)$ is defined only on the discrete frequency points for a finite $N$, it has, in general, nonzero tail probabilities, a phenomenon known as spectral leakage.
The QPE-based filtering amounts to discarding states with $y>y_{c}$ or $\omega_{y}>\omega_{y_{c}}$ in Eq. (3), where $y_{c}$ is a cutoff which is taken as a positive integer. To see the concept clearly, we rewrite the state in Eq. (3) as
$|\Psi\rangle$ $=$ $\sum_{y=0}^{y_{c}}\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}\Bigl[\frac{A_{\mu}(y)}{\sqrt{\mathcal{R}_{\mu}}}|y\rangle\Bigr]|\phi_{\mu}\rangle$ (8)
$+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle$
$=$ $\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}|\chi_{\mu}\rangle|\phi_{\mu}\rangle$
$+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle.$
Here, $\mathcal{R}_{\mu}$ is the renormalization factor, which depends on the state $|\phi_{\mu}\rangle$ through $P_{\mu}(y)$ in Eq. (5) as
$\mathcal{R}_{\mu}=\sum_{y=0}^{y_{c}}P_{\mu}(y),$ (9)
and
$|\chi_{\mu}\rangle=\sum_{y=0}^{y_{c}}\frac{A_{\mu}(y)}{\sqrt{\mathcal{R}_{\mu}}}|y\rangle$ (10)
are normalized $\mu$-dependent states. Thus, the QPE-based filtering amounts to extracting the first term of Eq. (8). One way to do this truncation is a coherent (i.e., fully quantum) filtering without measurement, by first performing an ideal quantum amplitude amplification (QAA) [gilles2002, PhysRevLett.113.210501, PRXQuantum.2.040203] and then tracing out the ancilla register. Another approach is through classical post-selection of the ancilla states after the measurement, as depicted in Fig. 1. In this case, the result becomes probabilistic, and the success probability can be increased by performing QAA on the corresponding states. This approach therefore differs from usual quantum filters but can be applied to enhance the spectral estimation, as discussed in Section III.
Figure 2: Errors in the QPE-based filtering. The filled regions indicate errors due to the truncation. (a) $P_{\mu}(y)$ for states with $E_{\mu}\leq\omega_{y_{c}}$. (b) $P_{\mu}(y)$ for states with $E_{\mu}>\omega_{y_{c}}$.
In both quantum and classical filtering approaches, the net effect of filtering is represented as the renormalization of the weight of each state $|\phi_{\mu}\rangle$ (see the Appendix A for additional explanation on the classical approach):
$|C_{\mu}|^{2}\to|C_{\mu}|^{2}\mathcal{R}_{\mu}.$ (11)
The ideal filtering corresponds to $\mathcal{R}_{\mu,\textrm{ideal}}=1$ for the states with $0\leq E_{\mu}\leq\omega_{y_{c}}$ and $\mathcal{R}_{\mu,\textrm{ideal}}=0$ otherwise. The deviation from this ideal filtering, denoted as $\Delta\mathcal{R}_{\mu}$, is given as
$\Delta\mathcal{R}_{\mu}=\Bigl\{\begin{array}[]{ll}\sum_{y=y_{c}+1}^{N-1}P_{\mu}(y)&E_{\mu}\leq\omega_{y_{c}}\\ \sum_{y=0}^{y_{c}}P_{\mu}(y)&E_{\mu}>\omega_{y_{c}},\end{array}\Bigr$. (14)
where $\sum_{y=0}^{N-1}P_{\mu}(y)=1$ is used. These errors originate from the tail part of the probability distribution due to spectral leakage, as schematically shown in Fig. 2. Here the errors are indicated by the filled regions in Figs. 2(a) and 2(b) for states with $E_{\mu}\leq\omega_{y_{c}}$ and $E_{\mu}>\omega_{y_{c}}$, respectively. Note that due to periodicity the QPE-based filtering has two boundaries, at $y=y_{c}$ and at $y=0$.
Although it is not an essential point of this approach, in this work we choose $y_{c}$ as $y_{c}=2^{n-m}-1$ with $m(<n)$ a positive integer. With this choice, $\omega_{y_{c}}=\frac{2\pi}{NT}y_{c}\approx\frac{2\pi}{T}\frac{1}{2^{m}}$, and we post-select or amplify the states whose first $m$ bits in the ancilla register are 0 (ex. $|{\bf{00}}1101\rangle$ when $m=2$). This filtering corresponds to a low-pass filter, but it is straightforward to consider filtering for general frequency intervals [Jensen_2021].
II.2 Window functions
Figure 3: (a) $a_{j}$ in Eq. (1) for the three windows. (b) $P_{\mu}(y)$ for the three windows and (c) $P_{\mu}(y)$ in the logarithmic scale. The vertical dotted lines in (b) and (c) indicate the position of $E_{\mu}$. Here $N=64$.
Figure 4: Error estimate in the Kaiser window-based filter defined in Eq. (23) for $n=6$ ($N=64$) and $n=10$ ($N=1024$). The horizontal dashed line indicates $\epsilon_{\textrm{max}}=10^{-7}$ used in Sec. II.5.
The performance of the QPE-based filtering depends crucially on the form of $a_{j}$. In this work, we consider three forms of $a_{j}$: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. Note that many other windows have been proposed in classical signal processing [prabhu], and a detailed exploration of their characteristics is left for future work.
The first window is the rectangular window characterized by
$a_{j}=\frac{1}{\sqrt{N}},$ (15)
which is used in the conventional QPE algorithm and corresponds to setting $U_{\textrm{in}}=\textrm{H}^{\otimes n}$ with H the Hadamard gate. The root mean squared error (RMSE) of the estimated phase achieved by this window function is proportional to $1/\sqrt{N}$ known as the standard quantum limit.
The second window is the sine window originally proposed in quantum metrology [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015]:
$a_{j}=\sqrt{\frac{2}{N}}\sin\frac{\pi}{N}j.$ (16)
In this case, the resulting RMSE is proportional to $1/N$ [4655455], which is known as the Heisenberg-limited scaling. The corresponding $U_{\textrm{in}}$ can be constructed in $O(n)$ complexity [PhysRevD.106.034503].
The third window is the Kaiser window [1163349, PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], which is introduced as an approximation of the prolate spheroidal window [6773659, https://doi.org/10.1002/sapm196544199, Imai_2009, patel2024optimalcoherentquantumphase]. We consider the following form:
$a_{j}\propto I_{0}\left(\pi\alpha\sqrt{1-\left(\frac{2j}{N}-1\right)^{2}}\right),$ (17)
where $I_{0}$ is the modified Bessel function of the order zero and $\alpha$ is a positive parameter. Note that our form is slightly different from other works [PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], in that our definition $j$ goes from 0 to N1N-1. The Kaiser window and the prolate spheroidal window are more difficult to implement than the first two, but there have been proposals [mcardle2022quantumstatepreparationcoherent, patel2024optimalcoherentquantumphase]. In particular, the method of Ref. [mcardle2022quantumstatepreparationcoherent] uses QSVT with $O(n\sqrt[4]{\pi\alpha+1}(\pi\alpha+\log\epsilon^{-1}))$ gate complexity within error $\epsilon$. In Appendix B, we review this approach and estimate the cost of constructing the Kaiser window.
We compare the results of all three windows in Fig. 3. In Fig. 3(a) we plot $a_{j}$ for the three windows. The Kaiser window has a Gaussian-like structure more localized than the sine window, and its spread is controlled by the parameter $\alpha$. In Figs. 3(b) and 3(c) we compare $P_{\mu}(y)$ defined in Eq. (5) for the three windows, for a specific example of $|\phi_{\mu}\rangle$ with $\frac{NT}{2\pi}E_{\mu}=10.2$ and $N=64$. The $P_{\mu}(y)$ in the rectangular window has the highest peak but has large leakage, and the sine window result shows a smaller tail probability [PhysRevA.110.022618]. For these two windows, the expression of $A_{\mu}(y)$ in Eq. (4) can easily be obtained; for the rectangular window
$A_{\mu}(y)$ $=$ $\frac{1}{N}\frac{1-e^{iE_{\mu}TN}}{1-e^{i\Theta_{\mu}(y)}},$ (18)
and in the case of the sine window
$A_{\mu}(y)$ $=$ $\frac{\sqrt{2}}{N}\frac{(1+e^{iE_{\mu}TN})e^{i\Theta_{\mu}(y)}\textrm{sin}\frac{\pi}{N}}{(1-e^{i(\Theta_{\mu}(y)+\frac{\pi}{N})})(1-e^{i(\Theta_{\mu}(y)-\frac{\pi}{N})})},$ (19)
where $\Theta_{\mu}(y)=E_{\mu}T-\frac{2\pi}{N}y$. Using the inequalities
$|1-e^{i\theta}|\leq 2$ (20)
and
$\frac{1}{|1-e^{i\theta}|}\leq\frac{\pi}{2|\theta\ \textrm{mod}[-\pi,\pi)|},$ (21)
it can be shown that $P_{\mu}(y)=|A_{\mu}(y)|^{2}$ decays as $O(y^{-2})$ and $O(y^{-4})$ for the rectangular and sine windows, respectively. This fact indicates that when these windows are used as a filter, the error defined in Eq. (14) also decays polynomially in $\max\bigl\{|E_{\mu}|^{-1},|\omega_{y_{c}}-E_{\mu}|^{-1}\bigr\}$ for $E_{\mu}\leq\omega_{y_{c}}$ or in $\textrm{max}\bigl\{|E_{\mu}-\omega_{y_{c}}|^{-1},|\frac{2\pi}{T}-E_{\mu}|^{-1}\bigr\}$ for $E_{\mu}>\omega_{y_{c}}$.
The properties of probability distribution for the Kaiser window have been investigated in detail [PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327]. As can be seen in Fig. 3(c), there is a clear separation between the high-probability region called mainlobe, which is confined in $\approx 2\alpha$ QPE frequency grid points near the peak center, and the rather flat region outside the mainlobe, called sidelobes. The sum of $P_{\mu}(y)$ in the sidelobes gives the amount of the leakage $\epsilon$, which is related to $\alpha$ in the leading order by [PRXQuantum.5.010319]
$\alpha=O\left(\log\frac{1}{\epsilon}\right).$ (22)
To assess the error in the Kaiser window-based filter, we plot in Fig. 4 the quantity
$\epsilon_{\textrm{max}}=\max_{E_{\mu}}\left[1-\sum_{y}{}^{{}^{\prime}}P_{\mu}(y)\right],$ (23)
where the prime indicates that the sum is over the $\lceil 2\alpha+1\rceil$ highest-probability points near the peak center $E_{\mu}$. This quantity is usually larger than $\Delta\mathcal{R}_{\mu}$ in Eq. (14) and provides an estimate of the maximum error in the Kaiser window-based filter for the states with $\textrm{min}\{E_{\mu},\omega_{y_{c}}-E_{\mu}\}>\omega_{\lceil 2\alpha\rceil}$ or $\textrm{min}\{E_{\mu}-\omega_{y_{c}},\frac{2\pi}{T}-E_{\mu}\}>\omega_{\lceil 2\alpha\rceil}$. The error decreases exponentially in $\alpha$ as expected, and the result has a very weak dependence on the number of frequency grid points $N$.
II.3 QPE-based filters for the three windows
Figure 5: (a) Weight-renormalization factor (filter) $\mathcal{R}_{\mu}$ (Eq. (9)) for the rectangular and the sine windows (top) and for the Kaiser window with $\alpha=3$ and 66 (bottom). (b) Deviations of $\mathcal{R}_{\mu}$ from the ideal filter, $\Delta\mathcal{R}_{\mu}$ given in Eq. (14), on the logarithmic scale for $E_{\mu}\leq 2$ (left) and $E_{\mu}\geq 5$ (right). The vertical dotted line in the left panel shows the position of $\omega_{y_{c}}$, and the filled squares indicate the positions of the QPE frequency grid points $\omega_{y}$.
Figure 5(a) shows $\mathcal{R}_{\mu}$ (given in Eq. (9)) as a function of $E_{\mu}$ for the three window functions described in the previous section. The parameters used are $T=1$, $n=6$ ($N=64$) and $y_{c}=15$, corresponding to extracting the states in approximately one fourth of the entire energy range. The rectangular-window-based filter exhibits an oscillating behavior known as the Gibbs phenomenon [Jerri1998], which originates from the discontinuity of the truncated frequency spectrum. In the case of the other two windows this oscillation is suppressed at this scale. As shown in the insets of Figs. 5(a) and 5(b), compared to the rectangular and sine window-based filters, the Kaiser window-based filters have a larger transitional region where $\mathcal{R}_{\mu}$ deviates from one and zero. This reflects the fact that the probability distribution of the Kaiser window has a broader mainlobe as shown in Figs. 3(b) and 3(c).
In Fig. 5(b), the errors of these filters, $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), are plotted together with the QPE frequency grid points (filled squares). The sine window result also shows oscillation, though of smaller magnitude than the rectangular window. By analyzing $P_{\mu}(y)$ [PhysRevA.110.022618], it can be shown that these two filters are exact (i.e. $\Delta\mathcal{R}_{\mu}=0$) for $E_{\mu}=\omega_{y^{\prime}}$ and $E_{\mu}=\omega_{y^{\prime}+1/2}$ for the rectangular and sine windows, respectively, with $y^{\prime}$ any integer.
The Kaiser window-based filter has the following properties:
$1-\epsilon$ $\leq$ $\mathcal{R}_{\mu},\quad\mbox{for}\quad E_{\mu}\in\left[\frac{\delta}{2},E_{\textrm{targ}}+\frac{\delta}{2}\right],$ (24)
$\mathcal{R}_{\mu}$ $\leq$ $\epsilon,\quad\mbox{for}\quad E_{\mu}\in\left[E_{\textrm{targ}}+\frac{3}{2}\delta,\frac{2\pi}{T}-\frac{\delta}{2}\right].$ (25)
Here $E_{\textrm{targ}}<\omega_{y_{c}}$ is the width of the target energy interval, and $\delta$ is the width of the transitional region near $E_{\mu}=0$, $\omega_{y_{c}}$, and $\omega_{N-1}$. The total width of the latter region is $2\delta$. By comparing the Kaiser window results and the QPE frequency grid points, $E_{\textrm{targ}}$ and $\delta$ can be written approximately as
$\delta$ $\approx$ $\omega_{\lceil 2\alpha\rceil}=\frac{2\pi}{NT}\lceil 2\alpha\rceil,$ (26)
$E_{\textrm{targ}}$ $\approx$ $\omega_{y_{c}}-\delta.$ (27)
From Eq. (26) and using the fact that $\alpha$ and $\epsilon$ are related as $\alpha=O(\log\epsilon^{-1})$, one can obtain
$N\approx\frac{2\pi}{\delta T}2\alpha=O(\delta^{-1}\log\epsilon^{-1}).$ (28)
The corresponding number of ancilla qubits is $O\Bigl(\log(\delta^{-1}\log\epsilon^{-1})\Bigr)$.
In practice, from a given set of ($\epsilon,\delta,E_{\textrm{targ}}$), the parameter $\alpha$, the number of ancilla qubits $n$, and the cutoff parameter $y_{c}$ in the Kaiser window-based filter can be determined as follows: One can choose $\alpha$ from the relation between $\alpha$ and $\epsilon$ shown in Fig. 4. The width of the translational region $\delta$ and $\alpha$ are related by $\delta\approx\frac{2\pi}{2^{n}}\alpha$, from which one can determine the number of qubits $n$. The cutoff parameter $y_{c}$ can be selected from $2^{n}$ QPE grid points.
II.4 The Gibbs phenomenon and the $\sigma$-factor
As seen in the previous section, the QPE-based filter $\mathcal{R}_{\mu}$ generally exhibits oscillation known as the Gibbs phenomenon [Jerri1998]. To analyze this effect, let us first rewrite $\mathcal{R}_{\mu}$ as
$\mathcal{R}_{\mu}=\sum_{j=-(N-1)}^{N-1}\tilde{g}_{y_{c}}(j)\sigma_{j}e^{-iE_{\mu}Tj},$ (29)
where
$\tilde{g}_{y_{c}}(j)=\frac{e^{i\frac{\pi}{N}y_{c}j}}{N}\frac{\textrm{sin}\frac{\pi(y_{c}+1)j}{N}}{\textrm{sin}\frac{\pi j}{N}}$ (30)
is the discrete Fourier transform of the discretized step function $g_{y_{c}}(y)$:
$g_{y_{c}}(y)$ $=$ $\Bigl\{\begin{array}[]{ll}1&\quad\mbox{for}\quad 0\leq y\leq y_{c}\\ 0&\quad\mbox{for}\quad y_{c}<y\leq N-1\end{array}\Bigr.$. (33)
For $N\gg j$, Eq. (30) becomes $\tilde{g}_{y_{c}}(j)\to e^{i\frac{z_{c}}{2}j}\frac{\textrm{sin}\frac{z_{c}}{2}j}{\pi j}$ with $z_{c}=\frac{2\pi}{N}y_{c}$, which is the Fourier series expansion coefficient of the following $2\pi$-periodic function $G_{z_{c}}(z)$:
$G_{z_{c}}(z)$ $=$ $\Bigl\{\begin{array}[]{ll}1&\quad\mbox{for}\quad 0\leq z\leq z_{c}\\ 0&\quad\mbox{for}\quad z_{c}<z\leq 2\pi\end{array}\Bigr$. (36)
Equation (29) has the same form as the truncated Fourier series expansion of the step function with a factor $\sigma_{j}$, which is constructed from the QPE window function $a_{j}$
$\sigma_{j}=\sum_{j^{\prime}=0}^{N-1-|j|}a_{|j|+j^{\prime}}a_{j^{\prime}}.$ (37)
In the theory of the Gibbs phenomenon [Jerri1998], this $\sigma$-factor is called a “filter” (not to be confused with filtering of states discussed in earlier sections) and serves as a smoothing factor that improves the convergence of the truncated Fourier series. Several filters have been proposed in the classical Fourier spectral analysis. Some examples are the Lanczos filter and the raised cosine filter [Jerri1998, Vandeven1991].
The explicit form of $\sigma_{j}$ for the rectangular window is
$\sigma_{j}=\frac{N-|j|}{N}$ (38)
and for the sine window
$\sigma_{j}=\frac{1}{N\sin\frac{\pi}{N}}\Bigl[\textrm{sin}\frac{\pi|j|}{N}\cos\frac{\pi}{N}+(N-|j|)\cos\frac{\pi|j|}{N}\sin\frac{\pi}{N}\Bigr].$ (39)
In the limit $N\to\infty$, Eqs. (38) and (39) become
$\sigma(x)=1-|x|$ (40)
and
$\sigma(x)=\frac{1}{\pi}\sin(\pi|x|)+(1-|x|)\cos(\pi x),$ (41)
respectively, with $x=j/N$.
In Fig. 6, we plot the above two forms as a function of $x=j/N$ as well as the Kaiser window result calculated numerically. In the rectangular window case, $\sigma_{j}$ has a triangular structure with a singularity at $j=0$. In the case of the sine and the Kaiser windows, $\sigma_{j}$ is more localized at $x=0$, therefore it filters out the Fourier expansion coefficients $\tilde{g}_{y_{c}}(j)$ in Eq. (29) for large $j$. This accelerates the convergence of the Fourier series expansion, and yields a smoother $\mathcal{R}_{\mu}$, as observed in Fig. 5. More detailed analysis of the $\sigma$-factor may offer a way to improve the window function, but is beyond the scope of the present work.
Figure 6: Comparison of the $\sigma$-factors defined in Eq. (37) as a function of $x=j/N$ for $N=64$.
II.5 Comparison with QETU
Figure 7: Quantum circuit for QETU. Here $U_{HT/2}=e^{iHT/2}$.
[…]
Our interest here is to compare the performance of the two filtering approaches in terms of the number of queries to $U_{HT}$, which has the most significant impact on the circuit depth. For this purpose, we perform a numerical simulation to estimate the number of queries in both approaches for a given set of parameters $E_{\textrm{targ}},\delta$, and $\epsilon$. Although QETU can produce a polynomial function satisfying Eqs. (24) and (25), here we consider a more simple and practical filtering function $F(E)$, which corresponds to $\mathcal{R}_{\mu}$ in Eq. (9), that satisfies the following conditions:
$|F(E)|$ $\leq$ $1\quad E\in[0,\frac{2\pi}{T}]$ (42)
$1-F(E)$ $\leq$ $\epsilon\quad E\in[0,E_{\textrm{targ}}]$ (43)
$|F(E)|$ $\leq$ $\epsilon\quad E\in[E_{\textrm{targ}}+2\delta,\frac{2\pi}{T}].$ (44)
This function has the same target region $E_{\textrm{targ}}$ and the transitional region $2\delta$ as in Eqs. (24) and (25).
Note that the filtering function $F(E)$ is applied to the squared amplitude for each state as $|C_{\mu}|^{2}\to F(E_{\mu})|C_{\mu}|^{2}$, rather than $C_{\mu}$ itself. We therefore construct the function $f(w)$ such that
$|f(w)|$ $\leq$ $1\quad w\in[0,1]$ (45)
$f(w)$ $=$ $c\quad w\in[w_{+},1]$ (46)
$f(w)$ $=$ $0\quad w\in[0,w_{-}]$ (47)
where
$c$ $=$ $\frac{1}{2}\Bigl[1+\sqrt{1-\epsilon}\Bigr],$ (48)
$w_{+}$ $=$ $\cos\frac{E_{\textrm{targ}}}{4}T,$ (49)
$w_{-}$ $=$ $\cos\frac{(E_{\textrm{targ}}+2\delta)}{4}T.$ (50)
This function is related to $F(E)$ satisfying the conditions Eqs. (42-44) as $F(E)=|f(\cos\frac{ET}{4})|^{2}$. This function is approximated by Chebyshev polynomials up to even degree $d$ as
$f(w)\approx f_{d}(w)=\sum_{l=0}^{d/2}c_{2l}T_{2l}(w),$ (51)
where $T_{2l}(w)$ and $c_{2l}$ are the $2l$-th Chebyshev polynomials and expansion coefficients, respectively. We numerically determine the smallest $d$ in which $F_{d}(E)=|f_{d}(\textrm{cos}\frac{ET}{4})|^{2}$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$.
We compare two approaches to determine $c_{2l}$: (1) the polynomial approximation of the erf function discussed by Low and Chuang [low2017hamiltoniansimulationuniformspectral] and (2) the convex optimization approach [PRXQuantum.3.040305].
The first approach uses the polynomial approximation of the erf function $\textrm{erf}(kx)(k>0)$ [low2017hamiltoniansimulationuniformspectral] that is also used in the previous QPE-QETU comparison in Ref. [PRXQuantum.5.040339]:
$p_{\textrm{erf},k,d+1}(x)$ $=$ $\frac{2ke^{-k^{2}/2}}{\sqrt{\pi}}\Bigl(I_{0}(k^{2}/2)x+$ (52)
$\sum_{j=1}^{d/2}I_{j}(k^{2}/2)(-1)^{j}$
$\bigl(\frac{T_{2j+1}(x)}{2j+1}-\frac{T_{2j-1}(x)}{2j-1}\bigr)\Bigr).$
The function $f_{d}(w)$ is constructed following the approximation of the rectangular function [low2017hamiltoniansimulationuniformspectral]
$f_{d}(w)$ $=$ $c\Bigl[1-\frac{1}{2}\Bigl(p_{\textrm{erf},2k,d+1}((w+w_{m})/2)$ (53)
$+p_{\textrm{erf},2k,d+1}((-w+w_{m})/2)\Bigr)\Bigr]$
$w_{m}$ $=$ $\frac{w_{+}+w_{-}}{2}$ (54)
$k$ $=$ $\frac{\sqrt{2}}{w_{+}-w_{-}}\log^{1/2}\bigl(\frac{2}{\pi\tilde{\epsilon}^{2}}\bigr)$ (55)
$\tilde{\epsilon}$ $=$ $1-\sqrt{1-\epsilon}.$ (56)
This $f_{d}(w)$ $\tilde{\epsilon}-$approximates $f(w)$ with $d=O(\frac{1}{w_{+}-w_{-}}\log\tilde{\epsilon}^{-1})=O(\frac{1}{\delta}\log\epsilon^{-1})$ [low2017hamiltoniansimulationuniformspectral].
In the second approach, the expansion coefficients $c_{2l}$ are calculated with the convex-optimization-based approach proposed in Ref. [PRXQuantum.3.040305] and using the qsppack library [PhysRevA.103.042419, Wang2022energylandscapeof, Dong2024infinitequantum]. We use $M$ grid points $w_{m=0,1,\dots,M-1}$ in $[0,1]$ and solve the following optimization problem
$\min_{\{c_{2l}\}}\max\Bigl[$ $\max_{w_{m}\in[w_{+},1]}|f_{d}(w_{m})-c|,$
$\max_{w_{m}\in[0,w_{-}]}|f_{d}(w_{m})|\Bigr]$
s.t. $|f_{d}(w_{m})|\leq 1\quad(m=0,1,\dots,M-1).$ (57)
In both approaches, we check whether the function $F_{d}(E)$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$ by combining a grid search with a fine grid of $100,000$ points and an optimization routine using the scipy.optimization module [2020SciPy-NMeth]. As an example of the calculated results, $f_{d}(w)$ and $F_{d}(E)$ obtained for $\epsilon=0.01$ are shown in Figs. 8(a) and 8(b).
Figure 8: Polynomial approximations of (a) $f(w)$ and (b) $F(E)=|f(\cos\frac{ET}{4})|^{2}$ for $\epsilon=0.01$ calculated with the prescription of Low and Chuang [low2017hamiltoniansimulationuniformspectral] with degree 1226 and via the convex optimization [PRXQuantum.3.040305] with degree 144. The parameters used are $E_{\textrm{targ}}=1.4430,2\delta=0.2556$, and $T=1$. The two vertical dotted lines indicate the positions of $w_{\pm}$ in (a) and $E_{\textrm{targ}}$ and $E_{\textrm{targ}}+2\delta$ in (b). The inset in (b) shows a magnified view.
Figure 9: Minimum number of queries to $U_{HT}$ in QETU for three sets of ($E_{\textrm{targ}},\delta$). The corresponding values in the QPE-based filters are also shown for comparison. The straight lines show the results of the linear regression. (a) $E_{\textrm{targ}}=1.0589$ and $2\delta=1.0232$. (b) $E_{\textrm{targ}}=1.4430$ and $2\delta=0.2556$. (c) $E_{\textrm{targ}}=1.5069$ and $2\delta=0.1279$. In (c) the results for $\epsilon=10^{-4}$ and $\epsilon=10^{-5}$ with $M=800$ are not calculated as there are less sampled points than the required degree $d$.
Figure 9 shows the calculated number of queries to $U_{HT}$, which is half of the minimum degree $d$, for three different sets of $(E_{\textrm{targ}},\delta)$. Here $E_{\textrm{targ}}$ and $\delta$ are obtained numerically from the results of the Kaiser window-based filter (Fig. 5 (b)) with $\alpha=3$ and $n=6,8,9$, using $y_{c}=2^{n-2}-1$ and $\epsilon=10^{-7}$ (see also Fig. 4). The number of sampled points $M$ in the second approach is a hyperparameter, and we compare three cases, $M=800,3200$ and $12800$. In the figures, $\epsilon$ is varied from $10^{-1}$ to $10^{-2}$ ($10^{-2}$ to $10^{-5}$) for the first (second) approach. The errors in the QPE-based filters with the three window functions are also shown for comparison.
It can be seen that the number of queries in the first approach is noticeably larger than that in the second approach; for $\epsilon=10^{-2}$, queries in the former approach are $\approx 9$ times larger than those in the latter. As a result, the QETU-based filter in the first approach requires more queries than the QPE-based filters. This aligns with the conclusions drawn in Ref. [PRXQuantum.5.040339] based on the rectangular window, which state that the QPE-based filtering is more efficient than QETU.
The number of required queries in the second approach, on the other hand, is smaller than those in the QPE-based filters with the rectangular and sine windows. It decreases as the number of sampled points $M$ increases, which can be understood by noting that with a small $M$ the optimized coefficients tend to overfit to the sampled data points. Interestingly, for the three cases studied, the QETU results in the second approach extrapolated to $\epsilon=10^{-7}$ are very close to the results of the Kaiser window-based filter (crosses). Although the required queries depend on the form of the QETU filter function used, this result can partly be understood by noting that the Kaiser window is an approximation of the prolate spheroidal window, which is designed to maximize the probability of the mainlobe of the probability distribution $P_{\mu}(y)$ [patel2024optimalcoherentquantumphase]. This construction is approximately equivalent to minimizing the error of the QPE-based filter $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), potentially explaining the similarities between the QETU-based filter obtained from the optimization (Eq. (57)) and the Kaiser window-based filter. We note that the present calculation compares two slightly different filter forms, but our comparison provides insight into the Kaiser window-based filter as an alternative to QETU and other QSP-based filtering methods.
We next discuss two situations in which the QPE-based filter could be advantageous. The first case is when we have prior knowledge of the input spectrum. For example, when we know the lower bound of the input spectrum $E_{\textrm{min}}$ satisfying $\delta<E_{\textrm{min}}$, we can reduce the error $\Delta\mathcal{R}_{\mu}$ for the states with $E_{\mu}\in[\frac{2\pi}{T}-\frac{\delta}{2},\frac{2\pi}{T}]$ (see Fig. 5(b)) by shifting the Hamiltonian and the cutoff frequency as $H\to H-\frac{\delta}{2}$ and $\omega_{y_{c}}\to\omega_{y_{c}}-\frac{\delta}{2}$, respectively. For this shifted system, $\mathcal{R}_{\mu}$ for the Kaiser window-based filter satisfies
$1-\epsilon$ $\leq$ $\mathcal{R}_{\mu}\quad\quad E_{\mu}^{\prime}\in\left[0,E_{\textrm{targ}}\right],$ (58)
$\mathcal{R}_{\mu}$ $\leq$ $\epsilon\quad\quad E_{\mu}^{\prime}\in\left[E_{\textrm{targ}}+\delta,\frac{2\pi}{T}\right],$ (59)
where $E_{\textrm{targ}}$ is defined in Eq. (27) and $E_{\mu}^{\prime}$ are shifted eigenenergies. These two conditions are identical to the QETU ones given in Eqs. (43) and (44) except that $2\delta$ in Eq. (44) is replaced by $\delta$. Since the number of queries in QETU scales as $O(\delta^{-1}\log\epsilon^{-1})$, this replacement approximately doubles the number of necessary queries in QETU. Therefore, in this case the Kaiser window-based filter is expected to be more efficient than the optimization-based QETU filter. A similar argument can be applied when the upper bound of the input spectrum $E_{\textrm{max}}$ is known. Furthermore, no shifting is required if both $E_{\textrm{min}}$ and $E_{\textrm{max}}$ are known and satisfy $E_{\textrm{min}}>\frac{\delta}{2}$ and $E_{\textrm{max}}<\frac{2\pi}{T}-\frac{\delta}{2}$, respectively.
The second case is when a numerical optimization procedure cannot be performed as, for example, the required accuracy is very high; in our calculation we have difficulty obtaining reliable minimum degrees for $\epsilon<10^{-5}$. For very small $\delta$ and $\epsilon$ one may have to rely on simpler non-optimization-based approaches such as Low and Chuang’s procedure. In such cases, the QPE-based filter with the Kaiser window can be a competitive alternative to QETU.
Note that the QPE-based approach requires $n=O\Bigl(\log\bigl[\delta^{-1}\log\epsilon^{-1}\bigr]\Bigr)$ ancilla qubits, and there are additional costs from the preparation of the Kaiser window as well as QFT. The gate complexity of QFT is $O(n^{2})$, and as discussed in Appendix B, the gate complexity of the QSVT-based approach in Ref. [mcardle2022quantumstatepreparationcoherent] is $O\Bigl(n\sqrt[4]{\pi\alpha+1}(\pi\alpha+\log\epsilon^{-1})\Bigr)=$ $O\Bigl(\log\Bigl[\delta^{-1}\log\epsilon^{-1}\Bigr](\log\epsilon^{-1})^{5/4}\Bigr)$ with three extra ancilla qubits. These costs are exponentially smaller than that of $N=2^{n}$ queries to $U_{HT}$ and $U_{HT}^{\dagger}$, therefore they do not become the bottleneck of the calculation for large systems. We also note that QETU requires $O(N)$ single-qubit rotations (see Fig. 7), and this cost is absent in the QPE-based filtering.
Finally, we briefly discuss general errors in practical calculations for both QPE and QETU approaches, which are ignored in the simulation. The most important error comes from time propagation. Using Trotterization [PhysRevX.11.011020], the time propagation operator $U_{HT}$ is approximated as $U_{HT}=\Bigl(\exp[iHT/r]\Bigr)^{r}\approx S_{p}(T/r)^{r}$, where $r$ is the Trotter step, and $S_{p}(t)$ is the $p$-th order Trotter product. The Trotter error can be evaluated in the operator norm
$U_{HT}-S_{p}(T/r)^{r}||=O(C_{\textrm{Trotter}}r^{-p}),$ (60)
where $C_{\textrm{Trotter}}$ is a system- and $p$-dependent constant factor constructed from nested commutators [PhysRevX.11.011020]. A number of works have investigated tighter, more practical bounds [10.5555/2871401.2871402, PhysRevA.91.022311, Su2021nearlytight, PhysRevA.105.012403, Sahinoglu2021, Chen2024, PhysRevLett.129.270502, blunt2025montecarloapproachbound]. With $N$ queries to $U_{HT}$ and $U_{HT}^{\dagger}$, the required steps for precision $\epsilon_{\textrm{Trotter}}$ is $r=O(N^{\frac{1}{p}}C_{\textrm{Trotter}}^{\frac{1}{p}}\epsilon_{\textrm{Trotter}}^{-\frac{1}{p}})$ [PRXQuantum.3.040305]. To avoid the Trotter error, qubitization and quantum walker based approaches have been proposed [Low2019hamiltonian, Berry2018], and they can also be combined with the current filtering.
In the early fault-tolerant era the imperfectness of quantum hardware also affects the results. Various quantum error mitigation techniques have to been proposed to remove errors for expectation value estimation and for sampling [RevModPhys.95.045005, liu2025quantumerrormitigationsampling], and these techniques can be applied to obtain correct outputs.
III Application to calculations of low-energy spectra
Figure 10: Quantum circuit for the QPE-based spectral simulation with filtering.
III.1 QPE-based spectral simulation
Finally, we consider an application of the present approach to the QPE-based calculation of the dynamical response function [PhysRevA.92.062318, PhysRevC.100.034610, PhysRevA.101.012330, PhysRevA.110.022618, PhysRevA.110.L060401, fomichev2024simulatingxrayabsorptionspectroscopy, nishi2025demonstrationlogicalquantumphase]. The objective is to calculate
$S_{B}(\omega)=\sum_{\mu}|\langle\phi_{\mu}|B|\phi_{0}\rangle|^{2}\delta(\omega-\Delta E_{\mu}),$ (61)
where $B$ is an operator which is not necessarily unitary, and $\Delta E_{\mu}=E_{\mu}-E_{0}$. The quantity $S_{B}(\omega)$ is related to the Fourier transform of the dynamical response function or the correlation function in the time domain and can be compared with various experimental spectra.
In the QPE-based approach, $S_{B}(\omega)$ is obtained using the quantum circuit in Fig. 1 by modifying the state preparation operation $U_{\psi}$ to block-encode $B|\phi_{0}\rangle$ with $a$ ancilla qubits as
$\Bigl[\langle 0|^{\otimes a}\otimes I\Bigr]U_{\psi}|0\rangle^{\otimes a}|0\rangle^{\otimes s}$ $=$ $B|\phi_{0}\rangle$ (62)
$=$ $\sum_{\mu}\langle\phi_{\mu}|B|\phi_{0}\rangle|\phi_{\mu}\rangle,$
and by replacing $H$ in $U_{HT}$ with $H-E_{0}$ if necessary. By collecting the measurement outputs in Fig. 1, the following approximation of $S_{B}(\omega)$ at discrete frequency points $\omega_{y}$ is obtained:
$\tilde{S}_{B}(\omega_{y})=\sum_{\mu}|\langle\phi_{\mu}|B|\phi_{0}\rangle|^{2}P_{\mu}(y).$ (63)
Figure 10 shows the quantum circuit with filtering applied prior to the QPE operation. Such filtering is useful in cases where one is interested in some specific small energy region, as in the case of virtual screening in materials science. The first step of this circuit performs filtering with $n$ ancilla qubits followed by QAA, as indicated by the gate labeled “Filter + QAA” in Fig. 10. Here we apply the proposed QPE-based filtering. The ancilla register can either be traced out or reused after measurement. The second step is to perform another QPE to obtain $\tilde{S}_{B}(\omega_{y})$ with $n^{\prime}$ ancilla qubits and time step $T^{\prime}$. The parameters $n^{\prime}$ and $T^{\prime}$ as well as the window function in $U_{\textrm{in}}^{\prime}$ used in the second step can be chosen independently to the ones used in the filtering step.
The final frequency resolution of this approach is $\Delta\omega^{\prime}=\frac{2\pi}{N^{\prime}T^{\prime}}$, with $N^{\prime}=2^{n^{\prime}}$. The parameters $n$ and $T$ used in the first filtering step can be smaller than those used in the second step, as in the coarse QPE approach proposed in Ref. [PRXQuantum.5.040339]. In a naive application of QAA without filtering, the total number of queries to $U_{HT}$ is $O(p_{0}^{-1/2}\beta N^{\prime})$, where $\beta=T^{\prime}/T$ and $p_{0}$ is the probability of measuring $\tilde{S}_{B}(\omega_{y})$ in the energy range of interest. With filtering, the total number of queries is reduced to $O(p_{0}^{-1/2}N+\beta N^{\prime})$, with $N=O(\delta^{-1}\log\epsilon^{-1})$ for the case of Kaiser window-based filter. This two-step approach therefore allows for the efficient calculation of the energy spectrum of the specified small energy region when $\beta N^{\prime}\gg N$ and can be useful to get fine structures of the spectrum at high resolution.
III.2 Density of states of MnO
Figure 11: Occupied part of the DOS of MnO (a) for the entire energy region and (b, c) near the cutoff after filtering using the rectangular window (b) and the Kaiser window with $\alpha=3$ (c). The solid curves show the DOS calculated with a fine QPE frequency grid, and the vertical dotted lines indicate the position of the cutoff frequency $\omega_{y_{c}}$. The states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$ are filtered out. The filled squares in (b) and (c) indicate the QPE frequency points used in filtering. Note that the Fermi level is not set to zero in this calculation.
Figure 12: (a) Error of the calculated DOS after filtering. (b) Contribution from the states with $\varepsilon_{\mathbf{k}\nu}\geq\omega_{y_{c}}$. (c) Contribution from the states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$. The vertical dotted lines indicate the position of $\omega_{y_{c}}$, and the filled squares in (b) and (c) indicate the QPE frequency points used in filtering.
To see the performance of the proposed QPE-based filtering for realistic spectra, we apply it to the calculation of the DOS [martin2020electronic] of the type-II antiferromagnetic MnO. This material is a prototype of transition metal monoxides, where strong electron correlation between the $3d$ orbitals of the transition metal atoms in a narrow energy region plays a crucial role in determining their electronic and magnetic properties. The insulating nature of the type-II antiferromagnetic MnO is explained by the combination of the magnetic ordering of Mn and the coupling between Mn $3d$ and O $2p$ orbitals [PhysRevB.30.4734, PhysRevLett.52.1830], and a number of band-structure calculations have been performed to understand the exact electronic structure of this system and also as a benchmark of a newly proposed algorithm  [PhysRevB.51.1477, PhysRevB.76.165106, PhysRevB.78.155112, PhysRevLett.103.036404, PhysRevB.79.235114, PhysRevX.5.011006, PhysRevB.101.165138].
In this work we focus on the occupied part of the DOS of this material, $D(\omega)$, within the one-particle approximation:
$D(\omega)=\sum_{\mathbf{k}\nu}^{\textrm{occupied}}\delta(\omega-\varepsilon_{\mathbf{k}\nu}),$ (64)
where $\varepsilon_{\mathbf{k}\nu}$ are one-particle orbital energies, and $\mathbf{k}$ and $\nu$ are the Bloch wavevectors and orbital (or band) indices [martin2020electronic], respectively. Although this DOS itself can be calculated efficiently on a classical computer, we use it as an approximation of the spectral function for interacting many-electron systems, which is hard to calculate classically [PhysRevA.101.012330, PhysRevA.110.L060401, PhysRevA.110.022618]. The DOS and spectral function are related to the experimental spectra obtained in photoemission spectroscopy.
In this work we prepare $\varepsilon_{\mathbf{k}\nu}=E_{\mu=(\mathbf{k}\nu)}$ classically using density functional theory [dreizler2012density] and set
$B=\sum_{\mathbf{k}\nu}\hat{c}_{\mathbf{k}\nu},$ (65)
where $\hat{c}_{\mathbf{k}\nu}$ are the electron annihilation operators. The DOS is calculated via Eq. (63) using a noiseless simulator. The input data is prepared using QuantumEspresso [QE-2017, QE-2009] with the pseudopotentials [martin2020electronic] from its pseudopotential database [QE-PPDB]. The rocksalt structure is assumed with lattice constant $a=4.43$ Å. The generalized gradient approximation (GGA) [PhysRevLett.77.3865] is used for the exchange and correlation functional, and the kinetic energy cutoff for wavefunctions used is $47$ Ry. The one-particle energies $\varepsilon_{\mathbf{k}\nu}$ are calculated with $32\times 32\times 32$ $\mathbf{k}$-points after obtaining the electron charge density self-consistently. Although the number of states in this calculation is prohibitively large for early fault-tolerant quantum computers, the resulting DOS yields a continuous spectrum that helps us to evaluate the efficacy of the proposed approach.
Figure 11(a) shows the DOS of this system without filtering, which is obtained with the sine window function for $U_{\textrm{in}}^{\prime}$. The sharp peaks below $\approx-0.2$ hartree from the Mn $3s$ and $3p$ orbitals and O $2s$ orbital play a minor role in determining the properties of this material. The broader spectra above $\approx+0.2$ hartree consist mainly of O $2p$ and Mn $3d$ orbitals. Our interest here is to obtain the DOS of the Mn $3d$ orbitals at high resolution. We therefore set a cutoff $\omega_{y_{c}}\approx 0.378$ as shown in the vertical dotted line in Fig. 11(a), and filter out the states below this threshold. In the calculation, we change the sign of $\varepsilon_{\mathbf{k}\nu}$ and apply a low-pass filter without modifying the current formalism.
The spectrum above the cutoff consists mainly of Mn $3d$ states. As can be seen in Fig. 11(b), due to the crystal field effect it splits into two main peaks [PhysRevLett.52.1830, PhysRevB.30.4734]: the lower-lying $t_{2g}$ peak and the higher-lying $e_{g}$ peak. The $e_{g}$ peak lies approximately from $0.42$ to $0.46$ hartree, and we set this energy region as the target energy range. We add a padding of 0.5 hartree above the $e_{g}$ states. Figures 11(b) and 11(c) show the DOS (obtained via Eq. (63)) after QPE-based filtering in open circles, calculated with the rectangular window (Fig. 11(b)) and with the Kaiser window using $\alpha=3$ (Fig. 11(c)). Note in both cases the sine window function is used in the second step. The parameters used are $n=n^{\prime}=6$ ($N=N^{\prime}=128$) and $T=2,T^{\prime}=20$, and the coarse QPE frequency grid points used in the filtering step are shown by the filled squares.
At this scale both the rectangular and the Kaiser windows efficiently filter out the states below the cutoff, which mainly consist of O $2p$ states, and both results yield higher-resolution spectra of the target ($e_{g}$) as well as $t_{2g}$ states, in good agreement with the original DOS. The Kaiser window result shows a better agreement with the original DOS without filtering, as expected.
Figure 12(a) shows the error of the calculated DOS after filtering for the three window functions. This error is obtained as the deviation from the unfiltered high-resolution DOS calculated with the states above $\omega_{y_{c}}$,
$\tilde{D}(\omega)=\sum_{\mathbf{k}\nu}^{\textrm{occupied},\varepsilon_{\mathbf{k}\nu}\geq\omega_{y_{c}}}P_{\mu=(\mathbf{k}\nu)}(y).$ (66)
The error is largest in the rectangular window result, and the error in the Kaiser window result decreases after $\pm\alpha$ frequency grid points used in filtering from the cutoff frequency $\omega_{y_{c}}$, in accordance with the observation in the previous section. Switching from the rectangular window to the Kaiser window reduces the error in the target energy region about $0.42$ to $0.46$ hartree, which corresponds to the $e_{g}$ orbital energies (see Figs. 11(b) and 11(c)), by about six orders of magnitude.
In Figs. 12(b) and 12(c) we decompose the error into the contribution from the states with (b) $\varepsilon_{\textbf{k}\nu}\geq\omega_{y_{c}}$ (i.e., Mn $t_{2g}$ and $e_{g}$) and (c) $\varepsilon_{\textbf{k}\nu}<\omega_{y_{c}}$. By comparing Figs. 12(a) and 12(b), it can be seen that the dominant error in the target region ($e_{g}$ states of Mn) originates from the states with $\varepsilon_{\textbf{k}\nu}\geq\omega_{y_{c}}$. Similarly, Fig. 12(c) shows the states with $\varepsilon_{\mathbf{k}\nu}<\omega_{y_{c}}$ are the main source of the error in region $\omega<\omega_{y_{c}}$. In Fig. 12(c) some extra low-intensity peaks are seen in the region $\omega\geq\omega_{y_{c}}$. These peaks originate from the low-lying states (see Fig. 11(a)) whose energies $\varepsilon_{\mathbf{k}\nu}$ are related to $\omega$ via $\omega=\varepsilon_{\mathbf{k}\nu}+\frac{2\pi}{T^{\prime}}r$ with $r$ a positive integer, as phases in QPE are defined modulo $2\pi$. This error is also mostly suppressed in the Kaiser window result.
In summary, this calculation demonstrates that the proposed two-step approach provides an efficient way of calculating high-resolution spectra of the specified small energy region. This approach is applicable to general discrete and continuous spectral simulations of quantum many-body systems, which are hard to calculate via classical means. Our results show that the simplest rectangular window-based filter already reproduces the main peak structures of the system; as the rectangular and sine windows are easy to construct, they can prove useful when the required accuracy is not very high. The Kaiser window-based filter provides an accurate and controllable approach, with tunable parameters $N$ and $\alpha$ based on the specified target energy interval and the tolerance error $\epsilon$ using Eqs. (26) and (27).
IV conclusions
In this work we have investigated the properties of the QPE-based filtering approach and have shown that the choice of window functions significantly affects the quality of the results. The results with the conventional rectangular window exhibit an oscillation known as the Gibbs oscillation, which originates from large spectral leakage for this window. This oscillation is suppressed when using the sine and the Kaiser windows, and this difference is explained by the $\sigma$-factor (smoothing factor) appearing in the analysis of the Gibbs oscillation. Further analysis and tuning of this factor may provide a way of improving window functions for QPE.
We have also compared the performance of the QPE-based filter and QETU in terms of necessary queries to the time propagation operator. We have found that, under the same conditions, the number of queries required in the Kaiser window-based filter is comparable to that in QETU after optimization of phase angles. This indicates that the form of the Kaiser window function is nearly optimal for filtering and can be used as a potential alternative to QETU. We have also estimated the cost of window preparation and have shown that it is exponentially smaller than that of the time propagation operations. We have also discussed that the Kaiser window-based filter can be more cost-effective than QETU when we have prior knowledge of the energy spectrum or filter construction in QETU is computationally hard.
As an application of the current approach, we have considered the combination of the filtering technique with the QPE-based spectral calculation. To assess the performance of the QPE-based filters for realistic spectra in materials science, we have performed a numerical simulation of the DOS of type-II antiferromagnetic MnO as an approximation of the spectral function. Our results have shown that the QPE-based filtering efficiently extracts the states in the energy range of interest, in this case the $e_{g}$ orbitals of Mn. Even the simplest rectangular window sufficiently reproduces the peak structures of this system, and the results can further be refined in a controllable manner by using the Kaiser window. These results indicate that, combined with efficient ground-state preparation and time evolution techniques, the proposed algorithm can be used as a simple QPE-based scheme complementary to QSP-based approaches for future materials simulations on a fault-tolerant quantum computer.
Acknowledgements.
We thank Hiroshi Yano, Jumpei Kato, Kohei Oshio, and Pawan Poudel for useful discussions. This work was supported by MEXT Quantum Leap Flagship Program Grants No. JPMXS0118067285 and No. JPMXS0120319794. K.W. was supported by JSPS KAKENHI Grant Number JP24KJ1963. K.S. acknowledges support from Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from JSPS, Japan.
Appendix A The case of classical post-selection
Here we provide a detailed explanation of the classical filter based on post-selection. Recall that the output state of the QPE circuit (Fig. 1) before the measurement is given by
$|\Psi\rangle$ $=$ $\sum_{y=0}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle$
$=$ $\sum_{\mu}C_{\mu}\sqrt{\mathcal{R}_{\mu}}|\chi_{\mu}\rangle|\phi_{\mu}\rangle+\sum_{y=y_{c}+1}^{N-1}\sum_{\mu}C_{\mu}A_{\mu}(y)|y\rangle|\phi_{\mu}\rangle.$
When the measurement outcome is $y$, the state after the measurement becomes
$\frac{\sum_{\mu}C_{\mu}A_{\mu}(y)|{\phi_{\mu}}\rangle}{\sqrt{\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y)|^{2}}},$ (67)
and the squared amplitude of state $|\phi_{\mu}\rangle$ is
$\frac{|C_{\mu}|^{2}|A_{\mu}(y)|^{2}}{\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y)|^{2}}.$ (68)
The expectation value of this squared amplitude is obtained by multiplying Eq. (68) by the conditional probability
$\frac{\sum_{\mu}|C_{\mu}|^{2}|A_{\mu}(y)|^{2}}{\sum_{y^{\prime}=0}^{y_{c}}\sum_{\mu^{\prime}}|C_{\mu^{\prime}}|^{2}|A_{\mu^{\prime}}(y^{\prime})|^{2}}$ (69)
and summing over $y=0,1,\dots,y_{c}$. This cancels out the denominator of Eq. (68) and yields $|C_{\mu}|^{2}\mathcal{R}_{\mu}$ up to a constant factor. Therefore Eq. (11) holds for the classical case as well as the quantum (coherent) case.
Appendix B Cost of constructing the Kaiser window via QSVT
In Ref. [mcardle2022quantumstatepreparationcoherent], a QSVT-based approach is presented to prepare a definite-parity function $f$ using QSVT and QAA with at most three ancilla qubits. This approach creates the operator
$U_{\textrm{QSVT+QAA}}|000\rangle|0\rangle^{\otimes n}=$
$\qquad\mathcal{N}\sum_{x=-\frac{N}{2}}^{\frac{N}{2}-1}\tilde{f}(\bar{x})|000\rangle|x\rangle,$ (70)
where $N=2^{n}$, $\bar{x}=\frac{2x}{N}$, $\mathcal{N}=\Bigl[\sum_{x=-\frac{N}{2}}^{\frac{N}{2}-1}|\tilde{f}(\bar{x})|^{2}\Bigr]^{-\frac{1}{2}}$, and $\tilde{f}(\bar{x})$ is an approximation of $\frac{f(\bar{x})}{\max_{\bar{x}\in[-1,1]}|f(\bar{x})|}$. Note that the right hand side of Eq. (70) can be converted to $|000\rangle\sum_{j=0}^{N-1}a_{j}|j\rangle$ used in our window state convention by operating a single $X$ gate.
Two building blocks are used to construct $U_{\textrm{QSVT+QAA}}$ as follows:
Figure 13: (a) Approximation of the Kaiser window $a_{j}^{d}$ with degree $d=4$ and $12$ for $\alpha=6$ and $n=8$. (b) Absolute error of the approximated window function $\Delta a_{j}=|a_{j}^{d}-a_{j}^{\textrm{exact}}|$ for $\alpha=6$ and $n=8$. (c) Total number of queries to $U_{\textrm{sin}},U_{\textrm{sin}}^{\dagger}$ in $U_{\textrm{QSVT+QAA}}$ (Eq. (70)) and the trace distance $\epsilon_{td}=\sqrt{1-(\sum_{j=0}^{N-1}a_{j}^{d}a_{j}^{\textrm{exact}})^{2}}$.
  1. 1.
    Block-encoding of the sine function, $U_{\sin}$, with one ancilla qubit
    $\Bigl(|0\rangle\langle 0|\otimes I\Bigr)U_{\sin}|0\rangle|x\rangle=\sin\bar{x}|0\rangle|x\rangle.$ (71)
    This operation requires $n+1$ single-qubit rotations and two multitarget CNOT gates.
  2. 2.
    QSVT operator $U_{\tilde{f}}$ that constructs $\tilde{f}(\bar{x})$ with another ancilla qubit
    $\Bigl(|00\rangle\langle 00|\otimes I\Bigr)U_{\tilde{f}}|00\rangle|x\rangle$ $=$ $\tilde{f}(\bar{x})|00\rangle|x\rangle$ (72)
    $=$ $h(\sin\bar{x})|00\rangle|x\rangle.$
    Here $h(y)$ is a degree-$d$ polynomial approximation of $\frac{f\bigl(\textrm{arcsin}(y)\bigr)}{\max_{\bar{x}\in[-1,1]}|f(\bar{x})|}$. This operation requires $d$ queries to $U_{\sin}$ and $U_{\sin}^{\dagger}$, $d$ single-qubit rotations, and $2d$ CNOT gates.
The operator $U_{\textrm{QSVT+QAA}}$ in Eq. (70) involves $2k+1$ operations of $U_{\tilde{f}}$ and $U_{\tilde{f}}^{\dagger}$, $2k+1$ single-qubit rotations, and $2k$ reflection operations. Here $k$ is the number of applications of the QAA operator, which is related to the L2 norm of the target state [mcardle2022quantumstatepreparationcoherent].
In the case of the Kaiser window
$f(\bar{x})=\frac{I_{0}(\pi\alpha\sqrt{1-\bar{x}^{2}})}{I_{0}(\pi\alpha)},$ (73)
the overall gate complexity of this approach is $O(n\sqrt[4]{\pi\alpha+1}(\pi\alpha+\log\epsilon_{td}^{-1}))$ [mcardle2022quantumstatepreparationcoherent], where $\epsilon_{td}$ is the error of the resulting window state in trace distance. Noting that the error in the Kaiser window-based filter $\epsilon$ and $\alpha$ are related by Eq. (22), by assuming $\epsilon\approx\epsilon_{td}$ we obtain $O(n(\log\epsilon^{-1})^{\frac{5}{4}})$ as the theoretical gate complexity of the Kaiser window construction.
To evaluate the actual cost of the window construction, we calculate optimal polynomial coefficients of the Kaiser window for a given degree $d$ and the corresponding phase angles using the qsppack library [PhysRevA.103.042419, Wang2022energylandscapeof, Dong2024infinitequantum] as done in Sec. II.5. In the calculation, we rescale the target function as $f\bigl(\textrm{arcsin}(y)\bigr)\to\eta f\bigl(\textrm{arcsin}(y)\bigr)$ with $\eta=0.98$ and use $500$ sample data points in $0,\sin(1)]$. This rescaling is to numerically ensure that $U_{\tilde{f}}$ is unitary and does not affect the normalization of the final window state [mcardle2022quantumstatepreparationcoherent]. In Fig. 13 (a) and Fig. 13 (b), we plot the approximations of the Kaiser window, $a_{j}^{d}$, and the corresponding error $|a_{j}^{d}-a_{j}^{\textrm{exact}}|$ with $d=4$ and $12$ for $\alpha=6$. Unlike the case with the step function discussed in Sec. II.5, we observe a fast convergence, as the Kaiser function has a well-behaved expansion.
In Fig. 13(c) we plot the trace distance between the approximated and exact window states $\epsilon_{td}=\sqrt{1-(\sum_{j=0}^{N-1}a_{j}^{d}a_{j}^{\textrm{exact}})^{2}}$ for $\alpha=3,6$, and $9$, as a function of the total number of queries to $U_{\textrm{sin}}$ and $U_{\textrm{sin}}^{\dagger}$. The total number of queries is given by the product of the degree $d$ and the number of QAA iterations $2k+1$, and the dominant gate cost of the calculation for large $n$ is obtained by multiplying this number by $n+1$ [mcardle2022quantumstatepreparationcoherent]. In the calculation $k$ is one for $\alpha=3$ and two for $\alpha=6$ and $9$, therefore QAA does not add a significant overhead. In the particular case of $\alpha=3$, the window state within error $\epsilon_{td}\approx 10^{-7}$, the same order as the error of the Kaiser window-based filter (see Fig. 4), can be obtained with $d=12$, corresponding to $36$ total queries to $U_{\textrm{sin}}$ and $U_{\textrm{sin}}^{\dagger}$. For $\alpha=6$ and $9$, the number of queries seems to depend sub-linearly on $\log\epsilon_{td}^{-1}$; our numerical fitting to $(\log\epsilon^{-1})^{\xi}$ gives $\xi\approx 0.5$.